### v1

# Scrutinize What We Ignore: Reining Task Representation Shift In Context-Based Offline Meta Reinforcement Learning

Abstract
Offline meta reinforcement learning (OMRL) has emerged as a promising approach for interaction avoidance and strong generalization performance by leveraging pre-collected data and meta-learning techniques. Previous context-based approaches predominantly rely on the intuition that maximizing the mutual information between the task and the task representation ($I(Z;M)$) can lead to performance improvements. Despite achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking. Motivated by the return discrepancy scheme in the model-based RL field, we find that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation. However, this optimization process ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse. To address this problem, we turn to use the framework of performance difference bound to consider the impacts of task representation shift explicitly. We demonstrate that by reining the task representation shift, it is possible to achieve monotonic performance improvements, thereby showcasing the advantage against previous approaches. To make it practical, we design an easy yet highly effective algorithm RETRO (REining Task Representation shift in context-based Offline meta reinforcement learning) with only adding one line of code compared to the backbone. Empirical results validate its state-of-the-art (SOTA) asymptotic performance, training stability and training-time consumption on MuJoCo and MetaWorld benchmarks.
1 Introduction
Reinforcement Learning (RL) has driven impressive advances in many complex decision-making problems in recent years [silver2018general, schrittwieser2020mastering], primarily through online RL methods. However, the extensive interactions of online RL entail high costs and safety concerns, posing significant challenges in real-world applications. Offline RL [wu2019behavior, levine2020offline] offers an appealing alternative by efficiently leveraging pre-collected data for policy learning, thereby circumventing the need for online interaction with the environment. This advantage extends the application of RL, covering healthcare [fatemi2022semi, tang2022leveraging], robotics [sinha2022s4rl, pmlr-v164-kumar22a] and games [schrittwieser2021online, mathieu2023alphastar]. Though demonstrating its superiority, offline RL holds a notable drawback towards generalizing to the unknown [ghosh2021generalization].
As a remedy, by combining the meta-learning techniques, OMRL [li2020focal, mitchell2021offline, xu2022prompting, li2024towards] has emerged as a beacon to illuminate the way toward strong generalization performance and fast adaptation capability while maintaining the merits of offline RL. Along the line of OMRL research, context-based OMRL (COMRL) algorithms [li2020focal, li2024towards] wield a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs). Specifically, they propose to train a context encoder independently to learn the task representation from the collected context (e.g. trajectories), then pass the latent representation to the downstream policy and value function, forming a decoupled framework. Thanks to [li2024towards] for unveiling the nature of the prior COMRL works that they predominantly rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvements [li2020focal, yuan2022robust, gao2024context, li2024towards]. Despite achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking.
Intriguingly, we unify the mutual information maximization objective commonly used in COMRL with the general RL objective of maximizing expected return. To be specific, maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation. This is achieved by extending the return discrepancy scheme [janner2019trust] to the COMRL framework. Thus, this analysis provides a feasible explanation for the performance improvement guarantee, filling the gap between intuition and theoretical justification. However, after scrutinizing this optimization process, we find it ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse. To address this problem, we adopt the framework of performance difference bound to consider the impacts of task representation shift explicitly. We theoretically demonstrate that by reining the task representation shift, it is possible to achieve monotonic performance improvements, thereby showcasing the advantage against prior works.
To make it practical, we derive an optimization problem concerning task representation shift and find that reining the update frequency of task representation and employing a straightforward classifier-based method for task representation extraction is sufficient to beat existing SOTA algorithms. RETRO is evaluated on a wide range of RL benchmarks covering MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta]. With the addition of just one line of code, RETRO achieves superior asymptotic performance and training stability compared to the classifier-based method alone. Notably, RETRO also holds a tuning-free property. Moreover, our approach to reining task representation shift is general and is also ablated in the experiments.
2 Related Works
Offline RL. Offline RL [fujimoto2019off] leverages a fully offline dataset to perform policy optimization and value function learning, avoiding explicit interaction with the environment. To overcome the notorious distribution shift and the induced value overestimation issues, prior works adopt methods based on policy regularization [fujimoto2019off, kumar2019stabilizing, fujimoto2021minimalist], value penalty [wu2019behavior, kumar2020conservative] and sequence modeling [chen2021decision, janner2021offline]. While these methods have demonstrated promising performance on a single task, their generalization performance to the unknown remains limited [ghosh2021generalization]. Our work follows the offline RL setting, attempting to adopt meta-learning techniques to improve the generalization performance.
Meta RL. Meta RL has emerged as an enticing avenue for effectively adapting to unknown environments [finn2017model]. Unlike traditional gradient-based meta RL methods [finn2017model, mitchell2021offline], context-based meta RL methods demonstrate superior advantages in both sample efficiency and asymptotic performance [rakelly2019efficient, zintgraf2019varibad, li2024towards]. Our work delves into this promising domain to advance the capabilities of context-based meta RL further.
Context-based Offline Meta RL. As the marriage between context-based meta RL and offline RL, COMRL combines the merits of both sides. Specifically, COMRL methods [li2020focal, yuan2022robust, gao2024context, li2024towards] leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input. At test time, COMRL methods leverage the generalization ability of learned task representation to perform meta-adaptation. Acknowledging the theoretical insights from [li2024towards], many prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] predominantly rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvements. Though achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking. We introduce the return discrepancy scheme [janner2019trust] into the COMRL setting to provide a feasible explanation. However, after scrutinizing this optimization process, we further find that it ignores the impacts of task representation shift, which may lead to performance improvement collapse. Centered around this motivation, our work aims to present a theoretical framework that reins the task representation shift.
Performance Improvement Guarantee. Performance improvement guarantee is a fundamental concern in both online RL and offline RL settings. In online RL, performance difference bound [schulman2015trust, kakade2002approximately, ji2022update, zhang2023how], return discrepancy [janner2019trust, luo2018algorithmic] and regret bound [osband2014model, curi2020efficient] are widely adopted schemes to get a performance improvement guarantee. In offline RL, CQL-based methods [kumar2020conservative, yu2021combo] also enjoy safe policy improvement guarantees. However, most works focus on the single-task setting, leaving the performance improvement guarantees in the multi-task and COMRL settings largely unexplored. We first provide a theoretical justification w.r.t the performance improvement guarantee for prior COMRL works. Then, we prove that it is possible to achieve monotonic performance improvements by reining the task representation shift ignored by prior COMRL works, thereby showcasing the advantage against them.
3 Preliminaries
3.1 Problem Statement
A task in RL is formalized as a fully observed MDP [puterman2014markov], which is represented by a tuple $(\mathcal{S},\mathcal{A},P,\rho_{0},R,\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(s^{\prime}|s,a)$, reward function $R(s,a)$, initial state distribution $\rho_{0}(s)$ and discount factor $\gamma\in[0,1]$. For arbitrary policy $\pi(a|s)$, we denote the state-action distribution at timestep $t$ as $d_{\pi,t}(s,a)\triangleq\text{Pr}(s_{t}=s,a_{t}=a|s_{0}\sim\rho_{0},a_{t}\sim\pi,s_{t+1}\sim P,\forall t\geq 0)$. The discounted state-action distribution of given $\pi(a|s)$ is denoted as $d_{\pi}(s,a)\triangleq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}d_{\pi,t}(s,a)$. The ultimate goal is to find an optimal policy $\pi(a|s)$ to maximize the expected return $\mathbb{E}_{(s,a)\sim d_{\pi}}[R(s,a)]$. A viable solution is to iterate the Bellman operator over the action-value function [sutton2018reinforcement], which is denoted as $Q_{\pi}(s,a)=\mathbb{E}_{\pi,P}[\sum^{\infty}_{t=0}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a]$.
It is widely assumed that the task $m$ in the meta RL setting is subject to the distribution $p(M)$ with sharing the same state space and action space but different from transition dynamics and reward functions [li2020focal, yuan2022robust, gao2024context, li2024towards]. At training time, tasks are randomly sampled from the task distribution $p(M)$. We denote the number of training tasks as $N$. For each task $i\in[0,1,...,N-1]$, an offline dataset $D_{i}=\{(s_{i,j},a_{i,j},s^{\prime}_{i,j},r_{i,j})\}_{j=1}^{K}$ is collected in advance, where $K$ denotes the number of transitions. The learning algorithm is required to train a meta-policy $\pi_{\text{meta}}$ with the only right to access the given offline datasets. At test time, given an unseen task, the meta policy $\pi_{\text{meta}}$ performs task adaptation to get a task-specific policy and then will be evaluated in the environment.
3.2 Context-based OMRL
COMRL holds a steadfast belief that the variation over different tasks can be represented by a compact representation $z$ [yuan2022robust]. In practice, COMRL methods would train a task encoder $Z(\cdot|x)$ to extract the task information [li2020focal, dorfman2021offline], where $x$ denotes the given context. Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x);\theta)$. Naturally, there exist the optimal task representation distribution $Z^{*}(\cdot|x)$ that can deduce the optimal expected return. The objective of COMRL is formalized as:
$\max_{\theta}J^{*}(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]$ (1)
where $R_{m}(s,a)$ denotes the ground-truth reward function of the task $m$.
Prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvement, as shown in Theorem 3.1.
Theorem 3.1 ([li2024towards]).
Denote $X_{b}$ and $X_{t}$ are the behavior-related $(s,a)$-component and task-related $(s^{\prime},r)$-component of the context $X$, with $X=(X_{b},X_{t})$. When meeting $I(Z;M;X_{b})\geq 0$:
$I(Z;X_{t}|X_{b})\leq I(Z;M)\leq I(Z;X)$ (2)
where 1) $\mathcal{L}_{FOCAL}\equiv-I(Z;X)=-I(Z;X_{t}|X_{b})-I(Z;X_{b})$; 2) $\mathcal{L}_{CORRO}\equiv-I(Z;X_{t}|X_{b})$; 3) $\mathcal{L}_{CSRO}\geq(\lambda-1)I(Z;X)-\lambda I(Z;X_{t}|X_{b})$
However, the theoretical justification of performance improvement for such intuition has been lacking.
4 Methods
To fill the gap, we introduce the return discrepancy scheme [janner2019trust] into the COMRL setting and find that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation (Section 4.1). Further, after scrutinizing this optimization process, we argue that it ignores the impacts of task representation shift, which may lead to performance improvement collapse. We then turn to use the framework of performance difference bound to consider task representation shift explicitly. We theoretically prove that it is possible to achieve monotonic performance improvements by reining task representation shift, thereby showcasing the advantage against prior works (Section 4.2). To make it practical, we derive an optimization problem concerning task representation shift and use the simplest algorithm to instantiate it (Section 4.3).
We first introduce the Lipschitz assumption for the following use. Lipschitz-based assumptions are widely adopted in both model-free RL [song2019efficient, ghosh2022provably] and model-based RL [ji2022update, zhang2023how].
Assumption 4.1.
The policy function is $L_{z}-$ Lipschitz with respect to some norm |||||$\cdot$||| | ⋅ | | in the sense that
$\forall Z_{1}(\cdot|x),Z_{2}(\cdot|x)\in\mathbb{Z},\,\,|\pi(\cdot|s,Z_{1}(\cdot|x);\theta)-\pi(\cdot|s,Z_{2}(\cdot|x);\theta)|\leq L_{z}\cdot|Z_{1}(\cdot|x)-Z_{2}(\cdot|x)|$ (3)
4.1 A Performance Improvement Perspective Towards Prior Works
Our goal is to induce an optimization objective served for COMRL to get the performance improvement guarantee. Motivated by the return discrepancy scheme [janner2019trust] in the model-based RL field, we can similarly build a tractable lower bound for $J^{*}(\theta)$.
Definition 4.2 (Return discrepancy in COMRL).
The return discrepancy in the COMRL setting can be defined as
$J^{*}(\theta)-J(\theta)\geq-|J(\theta)-J^{*}(\theta)|$ (4)
where $J(\theta)$ is the expected return of the policy $\pi$ among the expectation of tasks conditioning on the learned task embedding $Z(\cdot|x)$ as $J(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]]$.
From Definition 4.2, we can know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously. Then, Eq. (4) can be interpreted as a unified training objective for both the task encoder and the policy. Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage iteration framework with the first stage optimizing the task encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage optimizing the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$. Hence, the performance improvement guarantee can be achieved.
The following theorem induces a tractable upper bound for $|J^{*}(\theta)-J(\theta)|$, which forms a bedrock to connect the bridge between the intuition of prior works and the theoretical justification concerning the performance improvement guarantee.
Theorem 4.3 (Return bound in COMRL).
Denote $Z_{mutual}(\cdot|x)$ is the optimal task representation distribution served for maximizing $I(Z;M)$. Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
$|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (5)
Proof: See Section 7.2 in the Appendix. ∎
By directly unrolling the Theorem 4.3, we can establish the lower bound of $J^{*}(\theta)$ as:
$J^{*}(\theta)\geq J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (6)
Note that $|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$ is a constant w.r.t the learned task representation $Z(\cdot|x)$. Hence, this term can be ignored when optimizing $Z(\cdot|x)$. Admittedly, there may exist a smarter algorithm to further reduce the gap of $|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$, but this is not our main focus and we leave this to future work.
Theorem 4.3 and Eq. (6) indicate the following two principles for the task encoder and the policy. For the task encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$. For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$ as the task representation is irrelevant for policy optimization. To maximize $J(\theta)$, we can adopt standard offline RL algorithms [wu2019behavior, fujimoto2021minimalist, kumar2020conservative] similar to [yang2022unified].
Recall that prior COMRL works adopt alternatively training the context encoder by maximizing $I(Z;M)$ and training the policy network by conditioning on the learned task representation as well as applying the standard offline RL algorithms. As maximizing $I(Z;M)$ can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$, we argue that prior COMRL works can be interpreted by this optimization framework, thus filling the gap between the intuition and the theoretical justification.
However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, leaving the task representation shift between two consecutive updates alone. In the next section, we will show that the impacts of task representation shift may lead to performance improvement collapse.
4.2 Monotonic Performance Improvement Concerning Task Representation Shift
We attempt to build a theoretical framework that can achieve monotonic performance improvement. Hence, we turn to use the framework of performance difference bound.
Definition 4.4 (Performance difference bound in COMRL).
Denote $J^{1}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the task encoder and the policy. Similarly, denote $J^{2}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the task encoder and the policy. The performance difference bound in the COMRL setting can be defined as
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq C$ (7)
when $C$ is non-negative, the algorithm allows a monotonic performance improvement in the real environment.
According to Definition 4.4, we need to find the positive $C$ to improve the performance monotonically. To achieve this, we transform Eq. (7) to the following formulation, where $J^{mutual}(\theta_{i}),i\in[1,2]$ denote the expected return of the policy $\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{i})$ among the expectation of tasks.
$J^{2}(\theta_{2})-J^{1}(\theta_{1})$ $=(J^{2}(\theta_{2})-J^{mutual}(\theta_{2}))+(J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1}))$
$+(J^{mutual}(\theta_{1})-J^{2}(\theta_{1}))+(J^{2}(\theta_{1})-J^{1}(\theta_{1}))$ (8)
Then, we introduce a mild assumption for the task representation distribution.
Assumption 4.5.
The support set of $Z_{i}(\cdot|x),i\in[1,2]$ covers the support set of $Z_{mutual}(\cdot|x)$.
Assumption 4.5 usually holds in practice as when we give $z\sim Z_{mutual}(\cdot|x)$ to the task representation distribution $Z_{i}(\cdot|x)$ parameterized by the neural network, we can always get a non-zero probability. Given that augmented state $s^{+}=[s,z]$, by applying standard offline RL algorithms with enough samples, the policy update would lead to performance improvement when meeting Assumption 4.5, which means
$\epsilon^{mutual}_{12}\triangleq J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})>0$ (9)
By leveraging Eq. (4.2) and the above assumptions, we derive the lower bound of the performance difference.
Theorem 4.6 (Lower bound of performance difference in COMRL).
Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of performance difference in COMRL can be formalized as:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (10)
Proof: See Section 7.2 in the Appendix. ∎
For the need of monotonic performance improvement guarantee, we further let the right-hand side of Eq. (10) be positive. Then, we would have the following requirement to achieve monotonicity.
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (11)
The above equation unveils the key influence of the task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ for two aspects. If the task representation shift is too small, $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ may greater than the right-hand side of Eq. (11). On the contrary, if the task representation shift is too large, the right-hand side of Eq.(11) may be negative. Both of these cases would lead to performance improvement collapse. Thus, we argue that the impacts of task representation shift cannot be ignored.
To give an example for this problem, we introduce the following assumptions.
Assumption 4.7.
The impacts of task representation shift can be upper-bounded by $\beta$.
Assumption 4.8.
The space of the task representation is discrete and limited.
Assumption 4.9.
The i.i.d samples from $Z_{mutual}(\cdot|x)$ can be obtained.
Assumption 4.10.
The task encoder is trained by fitting an empirical distribution on i.i.d samples from $Z_{mutual}(\cdot|x)$.
Though sampling from $Z_{mutual}(\cdot|x)$ seems impractical, we can approximate it as sampling from a task encoder pre-trained by various bounds of $I(Z;M)$.
Corollary 4.11 (Monotonic performance improvement guarantee on empirical distribution).
Denote $\kappa$ as $\frac{2R_{max}}{(1-\gamma)^{2}}$. Given that the task encoder has already been trained by $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8, Assumption 4.9 and Assumption 4.10, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the task encoder from extra $k$ i.i.d samples, where:
$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (12)
Here, $\xi\in[0,1]$ is a constant.
Proof: See Section 7.2 in the Appendix. ∎
4.3 Practical Implementation
To simplest make it practical, we first transform Eq. (11) to an optimization problem.
$\min|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|,\,\,s.t.\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|>0$ (13)
As alluded in Section 4.1, many contrastive learning-based algorithms involved in prior works [li2020focal, yuan2022robust, gao2024context, li2024towards] can be adopted to minimize $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$. Unfortunately, no one among them is free from parameter-tuning. We propose to use a classifier-based manner to address this concern. Specifically, after extracting from the task encoder, the task representation would be passed through a fully connected layer to get the probabilities for each task. Then, the task encoder is trained by the supervision of the task label and back-propagation of the cross-entropy loss. Note that this classifier-based manner enjoys tuning-free property and also can be seen as an approximation of $I(Z;M)$. To the best of our knowledge, we are the first to propose adopting a classifier-based manner to train the task encoder in the COMRL setting.
For reining task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, since the values lying in the constraint of Eq. (13) are intractable, approximately estimating them may introduce a series of hyper-parameters thus significantly increasing the complexity of the algorithm. Instead, one insight from Corollary 4.11 is that tuning the update of the task encoder may bring better performance improvements. Through the experiment, we find that adjusting the update frequency of the task encoder to half of the origin is sufficient to beat the existing SOTA algorithms. Empirical results demonstrate that the previously ignored task representation shift does make a remarkable difference.
Our main algorithm is shown in Algorithm 1. Compared to the backbone, RETRO only adds one line of code, which is colored by red. To maintain the consistency of prior COMRL works, we also adopt BRAC [wu2019behavior] to train the policy. Moreover, combining prior works with our approach to reining task representation shift is also ablated in the experiments.
Algorithm 1 RETRO Framework
0:  Offline Datasets $\bm{X}=\{x^{i}\}_{i=1}^{N_{train}}$, training tasks $\bm{M}=\{m^{i}\}_{i=1}^{N_{train}}$, initialized learned policy $\pi_{\theta}$, Q function $Q_{\psi}$, context encoder $q_{\phi}$, task batch size $N_{tbatch}$
1:  while not done do
2:     Sample task $\{m^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{M}$ and the corresponding replay buffer =$\mathcal{R}=\{x^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{X}$
3:     for step in each iteration do
4:        Sample context $\{\bm{c}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
5:        Obtain task representations $\{\bm{z}^{j}\sim q_{\phi}(\bm{z}|\bm{c}^{j})\}_{j=1}^{N_{tbatch}}$
6:        // Rein the task representation shift
7:        if step % 2 == 0 then
8:           Compute $\mathcal{L}_{\text{RETRO}}=\textit{CrossEntropyLoss}(z^{j},m^{j})$
9:           Update $\phi$ to minimize $\mathcal{L}_{\text{RETRO}}$
10:        end if
11:        Detach the task embeddings $\{\bm{z}^{j}\}_{j=1}^{N_{tbatch}}$
12:        Sample training data $\{\bm{d}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
13:        Compute $\mathcal{L}_{\text{BRAC}}$ to update $\theta,\psi$
14:     end for
15:  end while
5 Experiment
Our experiments are conducted to answer 4 questions: 1) Compared to existing SOTA algorithms, can RETRO achieve more stable performance improvements and higher asymptotic performance? 2) Can RETRO be integrated into prior works like UNICORN and FOCAL? 3) How does reining the task representation shift with different frequencies affect RETRO? 4) Compared to existing SOTA algorithms, can RETRO achieve lower training-time consumption?
5.1 Environments Settings
We adopt MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta] benchmarks to evaluate the algorithms. We choose 4 competitive baselines to validate our superiority. The baselines cover UNICORN [li2024towards], which adopts reconstruction loss to balance the impacts of spurious correlation, CSRO [gao2024context], which adopts CLUB loss to upper-bound the impacts of the behavior policy, FOCAL [li2020focal], which directly applies distance metric learning to train the task encoder. and Classifier, which is our backbone without reining the task representation shift. The implementation details of all algorithms are shown in Appendix Section 7.3. Note that all these baselines can be categorized into the return discrepancy scheme stated in Section 4.1. Following the protocol of UNICORN, we randomly sample 20 training tasks and 20 testing tasks from the task distribution. We train the SAC [haarnoja2018soft] agent from scratch for each task and use the collected replay buffer as the offline dataset. During the meta-testing phase, we adopt a fully offline setting in a few-shot manner like [li2020focal] that randomly samples a trajectory from the dataset as the context to obtain the task representation and then passes it to the meta-policy to complete the testing in the true environments.
5.2 Baseline Comparisons
Figure 1: Testing returns of RETRO against baselines on six benchmarks. Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 1 illustrates the learning curves of RETRO against other baselines on 20 testing tasks for each benchmark. The results showcase that RETRO can unanimously achieve SOTA asymptotic performance across all benchmarks. Compared to the backbone Classifier, with adding only one line of code to rein the task representation shift, RETRO achieves remarkable performance improvements (e.g. approximately 20% relative performance improvements on Ant-Dir and 25% relative performance improvements on Dial-Turn). It is worth noting that RETRO shows comparably stable performance improvements in all the benchmarks especially on Point-Robot and Dial-Turn. Recall that RETRO only needs to adjust the update frequency to half of its origin, thus leading to a tuning-free algorithm. On the contrary, both UNICORN and CSRO are especially sensitive to the hyper-parameter settings. This phenomenon is particularly evident in the extremely challenging benchmark Humanoid-Dir, which holds the dimension of state space 376. We argue that without carefully parameter-tuning trials, UNICORN and CSRO cannot exhibit the desired performance. Admittedly, though more advanced algorithms may exist to better control the task representation shift, adopting the simplest manner to achieve SOTA performance makes RETRO more applicable to other fields.
5.3 Can RETRO Be Integrated Into Other Methods?
According to our analysis in Section 4.1, prior COMRL works ignore the impacts of task representation shift, which may lead to performance collapse. As our main focus is to rein the task representation shift, which is orthogonal to the prior works, RETRO should bring better performance improvements when being integrated into prior works. We denote RETRO-UNICORN and RETRO-FOCAL as combining the reining method into the original UNICORN and FOCAL respectively. To ensure fairness, we also adjust the update frequency of the task encoder in RETRO-UNICORN and RETRO-FOCAL to half of the origin.
Figure 2: Testing returns of RETRO-UNICORN and RETRO-FOCAL against the origin on three benchmarks. Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 2 demonstrates the learning curves of RETRO-UNICORN and RETRO-FOCAL against their backbones on 20 testing tasks for each benchmark. The results illustrate that adopting RETRO enables better performance improvements across all the benchmarks, thereby validating our core claim on reining the impacts of task representation shift. Note that RETRO(-Classifier) has comparable or even higher performance than RETRO-UNICORN, we believe that RETRO(-Classifier) should be a preferred option due to the tuning-free property.
Though our goal is to design the simplest algorithm to prove the effectiveness of reining the task representation shift, we can also notice that the convergence speed is relatively lower than the backbones at the beginning of the training process. Hence, in the future, we envision designing a strong algorithm to control the task representation shift delicately, where the springboard is RETRO.
5.4 The Influence Of Different Update Frequencies
Figure 3: Testing returns of difference update frequency on two benchmarks. Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 4: Average training time consumption of RETRO against other baselines on six benchmarks.
Figure 4 illustrates the influence of different update frequencies on the performance of RETRO. The results show that setting the update frequency to half of the origin can generally strike a balance among the convergence speed, training stability and asymptotic performance across all the benchmarks. Hence, we argue that adopting RETRO as a tuning-free algorithm is acceptable.
5.5 Can RETRO Reduce Training-Time Consumption?
We report the training-time consumption of RETRO against other baselines across all benchmarks in Figure 4. All algorithms are trained with the same architecture on two RTX3090 graphic cards. With the exception of Humanoid-Dir, which can only run one seed per graphic card due to the limitation, the reported times for the remaining benchmarks pertain to running two seeds per card and a total of four seeds concurrently during training. Figure 4 shows that RETRO has a significant time-consumption reduction against UNICORN [li2024towards], CSRO [gao2024context] and FOCAL [li2020focal]. Though this fact is based on our choice of backbone to be a classifier, it is the one proposed for the first time in our paper aiming at achieving a tuning-free algorithm. Compared to the backbone classifier, RETRO also achieves lower training-time consumption, as RETRO reduces the update frequency of the task encoder.
6 Conclusion & Limitation
Finding that the theoretical justification of maximizing $I(Z;M)$ can lead to performance improvement has been lacking, we constructively introduce the return discrepancy scheme into the COMRL setting, proving that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation. However, after scrutinizing this optimization framework, we find it ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse. To address this problem, we turn to use the framework of performance difference bound, proving that by reining the task representation shift, it is possible to achieve monotonic performance improvements. To make it practical, we propose a tuning-free algorithm RETRO with only adding one line of code compared to the backbone. Extensive experiments have validated the superiority of RETRO on training stability, asymptotic performance and training-time consumption. Though we choose classifier-based loss to train the task encoder, the idea that reins the task representation shift is general, which is shown in Section 5.3. We prospect that deeply exploring the role of task representation shift can make a profound difference in the COMRL setting. Naturally, one limitation is that there may exist more advanced algorithms to control the task representation shift, which is also stated in Section 5.3.
[…]
7 Appendix
7.1 Useful Lemmas
Lemma 7.1 (Return bound.[zhang2023how]).
Let $R_{max}$ denote the bound of the reward function, $\epsilon_{\pi}$ denote $\max_{s}D_{TV}(\pi_{1}||\pi_{2})$ and $\epsilon^{M_{2}}_{M_{1}}$ denote $\mathbb{E}_{(s,a)\sim d^{\pi_{1}}_{M_{1}}}[D_{TV}(p_{M_{1}}||p_{M_{2}})]$. For two arbitrary policies $\pi_{1},\pi_{2}\in\Pi$, the expected return under two arbitrary models $M_{1},M_{2}\in\mathcal{M}$ can be bounded as,
$|V^{\pi_{2}}_{M_{2}}-V^{\pi_{1}}_{M_{1}}|\leq 2R_{max}(\frac{\epsilon_{\pi}}{(1-\gamma)^{2}}+\frac{\gamma}{(1-\gamma)^{2}}\epsilon^{M_{2}}_{M_{1}})$ (14)
Lemma 7.2 (Inequility for $L_{1}$ deviation of the empirical distribution.[weissman2003inequalities]).
Let $P$ be a probability distribution on the set $\mathcal{A}=\{1,...,a\}$. For a sequence of samples $x_{1},...,x_{m}\sim P$, let $\hat{P}$ be the empirical probability distribution on $\mathcal{A}$ defined by $\hat{P}(j)=\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}(x_{i}=j)$. The $L_{1}$-deviation of the true distribution $P$ and the empirical distribution $\hat{P}$ over $\hat{A}$ from $m$ independent identically samples is bounded by,
$Pr(|P-\hat{P}|\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp{(-m\epsilon^{2}/2)}$ (15)
7.2 Missing Proofs
Theorem 7.3 (Return bound in COMRL).
Denote $Z_{mutual}(\cdot|x)$ is the optimal task representation distribution served for maximizing $I(Z;M)$. Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
$|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (16)
Proof:
$|J(\theta)-J^{*}(\theta)|$
$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]]-\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (17)
$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (18)
$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]$
$+\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (19)
$\leq\mathbb{E}_{m,x}[|\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]|$
$+|\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]|]$ (20)
$=\mathbb{E}_{m,x}[|\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)-\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)|$
$+|\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)-\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)|]$ (21)
$\leq\mathbb{E}_{m,x}[\frac{2R_{max}}{(1-\gamma)^{2}}\max_{s}D_{TV}(\pi(\cdot|s,Z(\cdot|x);\theta)||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta))$
$+\frac{2R_{max}}{(1-\gamma)^{2}}\max_{s}D_{TV}(\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)||\pi(\cdot|s,Z^{*}(\cdot|x);\theta))]$ (22)
$\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (23)
Eq. (22) is the result of directly applying Lemma 7.2 and Eq. (23) is the result of directly applying Assumption 4.1. ∎
Theorem 7.4 (Lower bound of performance difference in COMRL).
Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of performance difference in COMRL can be formalized as:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (24)
Proof:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})$ (25)
$=\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]$ (26)
$=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]]$ (27)
Denote $\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{i}(\cdot|x);\theta_{j})}}[R_{m}(s,a)]]$ as $G^{\theta_{j}}_{Z_{i}}$, then we can get the following derivation. Here, we only consider the value within the brackets.
$=\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]$ (28)
$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (29)
$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z_{mutual}}+G^{\theta_{2}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{mutual}}+G^{\theta_{1}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (30)
Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 7.1, Assumption 4.1 and Assumption 4.5 into the above formulation, we can get:
$G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z_{mutual}}+G^{\theta_{2}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{mutual}}+G^{\theta_{1}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (31)
$\geq-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{2}))+\epsilon^{mutual}_{12}$
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1}))$
$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1}))$ (32)
$\geq-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+\epsilon^{mutual}_{12}$
$-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (33)
Simplifying Eq. (33) we can get:
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (34)
Corollary 7.5 (Monotonic performance improvement guarantee on empirical distribution).
Denote $\kappa$ as $\frac{2R_{max}}{(1-\gamma)^{2}}$. Given that the task encoder has already been trained by $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$. When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8, Assumption 4.9 and Assumption 4.10, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the task encoder from extra $k$ i.i.d samples, where:
$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (35)
Here, $\xi\in[0,1]$ is a constant.
Proof: Let $Z_{1}$ has already been trained on $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$. Given that to get $Z_{2}$, we need to train $Z_{1}$ by extra $k$ i.i.d samples from $Z_{mutual}(\cdot|x)$ to get the monotonic performance improvement guarantee.
By applying Lemma 7.2, the $L_{1}$ deviation of the empirical distribution $Z_{2}(\cdot|x)$ and true $Z_{mutual}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
$Pr(|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|<=\epsilon)\geq 1-(2^{|\mathcal{\mathbb{Z}}|}-2)\exp{(-(N+k)\epsilon^{2}/2)}$ (36)
Then for a fixed $x$, with probability greater than $1-\xi$, we have:
$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{N+k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (37)
Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (38)
7.3 Implementation Details
In this section, we report the implementation details of RETRO and all the baselines.
  • FOCAL[li2020focal]. We adopt the open-source code of FOCAL [^1] https://github.com/FOCAL-ICLR/FOCAL-ICLR/. To better show the performance of FOCAL in each environment, we tune the parameter $\beta$ in distance metric learning from $[0.1,0.5,1]$ and choose the best result in our baseline comparison.
  • CSRO[gao2024context]. We adopt the open-source code of CSRO [^2] https://github.com/MoreanP/CSRO. To better show the performance of CSRO in each environment, we tune the club loss weight and task representation loss weight from $[10,20,30]\times[5,10,15]$ and choose the best result in our baseline comparison.
  • UNICORN[li2024towards]. We obtain the source code of UNICORN from the authors. To better show the performance of UNICORN in each environment, we tune the parameter $\frac{\alpha\beta-1}{(1-\alpha\beta)}$ from $[0.1,0.3,0.7,1,1.5]$ and choose the best result in our baseline comparison.
  • Classifier. Built upon the code base of FOCAL, Classifier only replaces the distance metric learning loss with cross-entropy loss, achieving tuning-free property.
  • RETRO. Built upon the code base of Classifier, RETRO only adjusts the update frequency of the task encoder to half of the origin, by adding merely one row of code.
To maintain fairness, we make sure all benchmark-irrelevant parameters are consistent, as shown in Table 1.
Table 1: Benchmark-irrelevant parameters setting in the training process.
training tasks 20
testing tasks 20
task training batch size 16
rl batch size 256
context size 1 trajectory
actor-network size [256, 256]
critic-network size [256, 256]
task encoder network size [64, 64]
learning rate 4e-3
Table 2: Benchmark-relevant parameters setting in the training process.
Configurations Point-Robot Humanoid-Dir Reach Ant-Dir Button-Press Dial-Turn
dataset size 2100 9e4 3e5 9e4 3e5 3e5
task representation dimension 5 64 5 5 5 5
For each environment, we make the benchmark-relevant parameters the same for each algorithm, as shown in Table 2.

### v2

@@
-Scrutinize What We Ignore: Reining Task Representation Shift In Context-Based Offline Meta Reinforcement Learning
+Scrutinize What We Ignore: Reining In Task Representation Shift Of Context-Based Offline Meta Reinforcement Learning
 Abstract
 Offline meta reinforcement learning (OMRL) has emerged as a promising approach for interaction avoidance and strong generalization performance by leveraging pre-collected data and meta-learning techniques.
-Previous context-based approaches predominantly rely on the intuition that maximizing the mutual information between the task and the task representation ($I(Z;M)$) can lead to performance improvements.
-Despite achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking.
-Motivated by the return discrepancy scheme in the model-based RL field, we find that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation.
-However, this optimization process ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse.
-To address this problem, we turn to use the framework of performance difference bound to consider the impacts of task representation shift explicitly.
-We demonstrate that by reining the task representation shift, it is possible to achieve monotonic performance improvements, thereby showcasing the advantage against previous approaches.
-To make it practical, we design an easy yet highly effective algorithm RETRO (REining Task Representation shift in context-based Offline meta reinforcement learning) with only adding one line of code compared to the backbone.
-Empirical results validate its state-of-the-art (SOTA) asymptotic performance, training stability and training-time consumption on MuJoCo and MetaWorld benchmarks.
+Previous context-based approaches predominantly rely on the intuition that alternating optimization between the context encoder and the policy can lead to performance improvements, as long as the context encoder follows the principle of maximizing the mutual information between the task and the task representation ($I(Z;M)$) while the policy adopts the standard offline reinforcement learning (RL) algorithms conditioning on the learned task representation.
+Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored.
+Inspired by the return discrepancy scheme in the model-based RL field, we find that the previous optimization framework can be linked with the general RL objective of maximizing the expected return, thereby providing a feasible explanation concerning performance improvements.
+Furthermore, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating optimization process, which may lead to performance improvement collapse.
+We name this issue task representation shift and theoretically prove that the monotonic performance improvements can be guaranteed with appropriate context encoder updates.
+We set different manners to rein in the task representation shift on three widely adopted training objectives concerning maximizing $I(Z;M)$ across different data qualities.
+Empirical results show that reining in the task representation shift can indeed improve performance.
+Our work opens up a new avenue for OMRL, leading to a better understanding between the performance and the task representation.
 1 Introduction
-Reinforcement Learning (RL) has driven impressive advances in many complex decision-making problems in recent years [silver2018general, schrittwieser2020mastering], primarily through online RL methods.
-However, the extensive interactions of online RL entail high costs and safety concerns, posing significant challenges in real-world applications.
+RL has driven impressive advances in many complex decision-making problems in recent years [silver2018general, schrittwieser2020mastering, zhang2023safe], primarily through online RL methods.
+However, the extensive interactions required by online RL entail high costs and safety concerns, posing significant challenges for real-world applications.
 Offline RL [wu2019behavior, levine2020offline] offers an appealing alternative by efficiently leveraging pre-collected data for policy learning, thereby circumventing the need for online interaction with the environment.
@@
 Though demonstrating its superiority, offline RL holds a notable drawback towards generalizing to the unknown [ghosh2021generalization].
-As a remedy, by combining the meta-learning techniques, OMRL [li2020focal, mitchell2021offline, xu2022prompting, li2024towards] has emerged as a beacon to illuminate the way toward strong generalization performance and fast adaptation capability while maintaining the merits of offline RL.
-Along the line of OMRL research, context-based OMRL (COMRL) algorithms [li2020focal, li2024towards] wield a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs).
-Specifically, they propose to train a context encoder independently to learn the task representation from the collected context (e.g.
-trajectories), then pass the latent representation to the downstream policy and value function, forming a decoupled framework.
-Thanks to [li2024towards] for unveiling the nature of the prior COMRL works that they predominantly rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvements [li2020focal, yuan2022robust, gao2024context, li2024towards].
-Despite achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking.
-Intriguingly, we unify the mutual information maximization objective commonly used in COMRL with the general RL objective of maximizing expected return.
-To be specific, maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation.
+As a remedy, by combining the meta-learning techniques, OMRL [li2020focal, mitchell2021offline, xu2022prompting] has emerged as an effective training scheme toward strong generalization performance and fast adaptation capability while maintaining the merits of offline RL.
+Among the OMRL research, context-based OMRL (COMRL) algorithms [li2020focal, li2024towards] hold a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs).
+Specifically, these methods [li2020focal, gao2024context, li2024towards] propose to train a context encoder via maximizing $I(Z;M)$ to learn the task representation from the collected context (e.g.
+trajectories) and train the downstream policy with standard offline RL algorithms conditioning on the task representation, as shown in Figure 1 left.
+The context encoder and policy are trained in an alternating manner, with each being updated once per cycle.
+Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored.
+Figure 1: Our training framework compared to the previous training framework.
+They both adopt the alternating optimization framework to train the context encoder and the policy.
+However, ours considers the variation of task representation ignored by the previous training framework.
+Intriguingly, we formalize this training framework as an alternating two-stage optimization framework and then link it with the general RL objective of maximizing the expected return.
+To be specific, maximizing $I(Z;M)$ and adopting standard offline RL algorithms can be interpreted as consistently raising the lower bound of the expected return conditioning on the optimal task representation distribution among the expectation of tasks.
 This is achieved by extending the return discrepancy scheme [janner2019trust] to the COMRL framework.
-Thus, this analysis provides a feasible explanation for the performance improvement guarantee, filling the gap between intuition and theoretical justification.
-However, after scrutinizing this optimization process, we find it ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse.
-To address this problem, we adopt the framework of performance difference bound to consider the impacts of task representation shift explicitly.
-We theoretically demonstrate that by reining the task representation shift, it is possible to achieve monotonic performance improvements, thereby showcasing the advantage against prior works.
-To make it practical, we derive an optimization problem concerning task representation shift and find that reining the update frequency of task representation and employing a straightforward classifier-based method for task representation extraction is sufficient to beat existing SOTA algorithms.
-RETRO is evaluated on a wide range of RL benchmarks covering MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta].
-With the addition of just one line of code, RETRO achieves superior asymptotic performance and training stability compared to the classifier-based method alone.
-Notably, RETRO also holds a tuning-free property.
-Moreover, our approach to reining task representation shift is general and is also ablated in the experiments.
+Thus, this analysis provides a feasible explanation for the performance improvement guarantee.
+More importantly, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating process.
+Variations of the task representation that are either too large or too small can lead to performance improvement collapse.
+To highlight the characteristic, we name this issue task representation shift and theoretically prove that it is possible to achieve monotonic performance improvement guarantees with proper updates.
+To show the impacts of this issue, we set different manners to rein in the task representation shift on contrastive objective adopted in [li2020focal, gao2024context, li2024towards] and reconstruction objective adopted in [zintgraf2019varibad, dorfman2021offline, li2024towards] across different data qualities.
+These two objectives are proven to maximize the approximate bounds of $I(Z;M)$ [li2024towards].
+We also introduce the cross-entropy objective, which is a straightforward discrete approximation to $I(Z;M)$.
+Empirical results show that reining in the task representation shift can indeed improve performance.
+Our work opens up a new avenue for OMRL, leading to a better understanding between the performance and the task representation.
 2 Related Works
@@
 Specifically, COMRL methods [li2020focal, yuan2022robust, gao2024context, li2024towards] leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input.
-At test time, COMRL methods leverage the generalization ability of learned task representation to perform meta-adaptation.
-Acknowledging the theoretical insights from [li2024towards], many prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] predominantly rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvements.
-Though achieving attractive results, the theoretical justification of performance improvement for such intuition has been lacking.
-We introduce the return discrepancy scheme [janner2019trust] into the COMRL setting to provide a feasible explanation.
-However, after scrutinizing this optimization process, we further find that it ignores the impacts of task representation shift, which may lead to performance improvement collapse.
-Centered around this motivation, our work aims to present a theoretical framework that reins the task representation shift.
+At test time, COMRL methods leverage the generalization ability of the learned context encoder to perform meta-adaptation.
+Acknowledging the theoretical insights from [li2024towards], prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] mainly focus on how to design a context encoder learning algorithm to better approximate $I(Z;M)$.
+However, our work moves the focus on how to get better asymptotic performance based on a given context encoder learning algorithm.
+Center around this motivation, we identify a new issue called task representation shift, which is ignored by the previous COMRL endeavors.
 Performance Improvement Guarantee.
-Performance improvement guarantee is a fundamental concern in both online RL and offline RL settings.
-In online RL, performance difference bound [schulman2015trust, kakade2002approximately, ji2022update, zhang2023how], return discrepancy [janner2019trust, luo2018algorithmic] and regret bound [osband2014model, curi2020efficient] are widely adopted schemes to get a performance improvement guarantee.
+Ensuring performance improvement is a key concern in both online and offline reinforcement learning settings.
+In online RL, performance improvement guarantees are often established through methods such as performance difference bounds [schulman2015trust, kakade2002approximately, ji2022update, zhang2023how], return discrepancy [janner2019trust, luo2018algorithmic], and regret bounds [osband2014model, curi2020efficient].
 In offline RL, CQL-based methods [kumar2020conservative, yu2021combo] also enjoy safe policy improvement guarantees.
-However, most works focus on the single-task setting, leaving the performance improvement guarantees in the multi-task and COMRL settings largely unexplored.
-We first provide a theoretical justification w.r.t the performance improvement guarantee for prior COMRL works.
-Then, we prove that it is possible to achieve monotonic performance improvements by reining the task representation shift ignored by prior COMRL works, thereby showcasing the advantage against them.
+However, most works focus on the single-task setting, leaving the performance improvement guarantees in the COMRL settings largely unexplored.
+While ContrBAR [choshen2023contrabar] also benefits from the performance improvement guarantee in the OMRL setting, it is specifically served for its own framework.
+In contrast, our work addresses a broader class of algorithms that maximize $I(Z;M)$, including but not limited to ContrBAR.
 3 Preliminaries
 3.1 Problem Statement
-A task in RL is formalized as a fully observed MDP [puterman2014markov], which is represented by a tuple $(\mathcal{S},\mathcal{A},P,\rho_{0},R,\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(s^{\prime}|s,a)$, reward function $R(s,a)$, initial state distribution $\rho_{0}(s)$ and discount factor $\gamma\in[0,1]$.
+A task in RL is generally formalized as a fully observed MDP [puterman2014markov], which is represented by a tuple $(\mathcal{S},\mathcal{A},P,\rho_{0},R,\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(s^{\prime}|s,a)$, reward function $R(s,a)$, initial state distribution $\rho_{0}(s)$ and discount factor $\gamma\in[0,1]$.
 For arbitrary policy $\pi(a|s)$, we denote the state-action distribution at timestep $t$ as $d_{\pi,t}(s,a)\triangleq\text{Pr}(s_{t}=s,a_{t}=a|s_{0}\sim\rho_{0},a_{t}\sim\pi,s_{t+1}\sim P,\forall t\geq 0)$.
@@
 The ultimate goal is to find an optimal policy $\pi(a|s)$ to maximize the expected return $\mathbb{E}_{(s,a)\sim d_{\pi}}[R(s,a)]$.
-A viable solution is to iterate the Bellman operator over the action-value function [sutton2018reinforcement], which is denoted as $Q_{\pi}(s,a)=\mathbb{E}_{\pi,P}[\sum^{\infty}_{t=0}\gamma^{t}r(s_{t},a_{t})|s_{0}=s,a_{0}=a]$.
-It is widely assumed that the task $m$ in the meta RL setting is subject to the distribution $p(M)$ with sharing the same state space and action space but different from transition dynamics and reward functions [li2020focal, yuan2022robust, gao2024context, li2024towards].
-At training time, tasks are randomly sampled from the task distribution $p(M)$.
+Given a distribution of tasks $p(M)$, every task $m$ sampled from $p(M)$ is an MDP sharing the same state space and action space but differing from transition dynamics and reward functions.
+It is widely assumed that the tasks in the meta RL setting are randomly sampled from $p(M)$ [li2020focal, yuan2022robust, gao2024context, li2024towards].
 We denote the number of training tasks as $N$.
 For each task $i\in[0,1,...,N-1]$, an offline dataset $D_{i}=\{(s_{i,j},a_{i,j},s^{\prime}_{i,j},r_{i,j})\}_{j=1}^{K}$ is collected in advance, where $K$ denotes the number of transitions.
-The learning algorithm is required to train a meta-policy $\pi_{\text{meta}}$ with the only right to access the given offline datasets.
+The learning algorithm is required to train a meta-policy $\pi_{\text{meta}}$ with access only to the given offline datasets.
 At test time, given an unseen task, the meta policy $\pi_{\text{meta}}$ performs task adaptation to get a task-specific policy and then will be evaluated in the environment.
 3.2 Context-based OMRL
-COMRL holds a steadfast belief that the variation over different tasks can be represented by a compact representation $z$ [yuan2022robust].
-In practice, COMRL methods would train a task encoder $Z(\cdot|x)$ to extract the task information [li2020focal, dorfman2021offline], where $x$ denotes the given context.
+Previous COMRL endeavors hold that the variation over different tasks can be represented by a compact representation $z$ [yuan2022robust].
+In practice, they choose to train a context encoder $Z(\cdot|x)$ to extract the task information [li2020focal, dorfman2021offline], where $x$ denotes the given context.
 Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x);\theta)$.
-Naturally, there exist the optimal task representation distribution $Z^{*}(\cdot|x)$ that can deduce the optimal expected return.
-The objective of COMRL is formalized as:
+Naturally, there exists an optimal task representation distribution $Z^{*}(\cdot|x)$ that meets the optimal return for any policy parameter $\theta$ is determined by $\pi(a|s,Z^{*}(\cdot|x);\theta)$.
+Hence, the objective of COMRL is formalized as:
 $\max_{\theta}J^{*}(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]$ (1)
 where $R_{m}(s,a)$ denotes the ground-truth reward function of the task $m$.
-Prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to rely on the intuition that maximizing $I(Z;M)$ can lead to performance improvement, as shown in Theorem 3.1.
+Previous COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to optimize the context encoder by maximizing the approximate bounds of $I(Z;M)$, as shown in Theorem 3.1.
+They hold the intuition that using such kind of manner to optimize the context encoder and adopting the standard offline RL algorithms to optimize the policy can lead to performance improvements.
+However, the theoretical justification of performance improvement for such intuition has been less explored.
 Theorem 3.1 ([li2024towards]).
-Denote $X_{b}$ and $X_{t}$ are the behavior-related $(s,a)$-component and task-related $(s^{\prime},r)$-component of the context $X$, with $X=(X_{b},X_{t})$. When meeting $I(Z;M;X_{b})\geq 0$:
+Denote $X_{b}$ and $X_{t}$ are the behavior-related $(s,a)$-component and task-related $(s^{\prime},r)$-component of the context $X$, with $X=(X_{b},X_{t})$. We have:
 $I(Z;X_{t}|X_{b})\leq I(Z;M)\leq I(Z;X)$ (2)
 where 1) $\mathcal{L}_{FOCAL}\equiv-I(Z;X)=-I(Z;X_{t}|X_{b})-I(Z;X_{b})$; 2) $\mathcal{L}_{CORRO}\equiv-I(Z;X_{t}|X_{b})$; 3) $\mathcal{L}_{CSRO}\geq(\lambda-1)I(Z;X)-\lambda I(Z;X_{t}|X_{b})$
-However, the theoretical justification of performance improvement for such intuition has been lacking.
 4 Methods
-To fill the gap, we introduce the return discrepancy scheme [janner2019trust] into the COMRL setting and find that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation (Section 4.1).
-Further, after scrutinizing this optimization process, we argue that it ignores the impacts of task representation shift, which may lead to performance improvement collapse.
-We then turn to use the framework of performance difference bound to consider task representation shift explicitly.
-We theoretically prove that it is possible to achieve monotonic performance improvements by reining task representation shift, thereby showcasing the advantage against prior works (Section 4.2).
-To make it practical, we derive an optimization problem concerning task representation shift and use the simplest algorithm to instantiate it (Section 4.3).
-We first introduce the Lipschitz assumption for the following use.
-Lipschitz-based assumptions are widely adopted in both model-free RL [song2019efficient, ghosh2022provably] and model-based RL [ji2022update, zhang2023how].
+We first introduce a Lipschitz assumption, a widely used technique in both model-free [song2019efficient, ghosh2022provably] and model-based [ji2022update, zhang2023how] RL frameworks.
 Assumption 4.1.
-The policy function is $L_{z}-$ Lipschitz with respect to some norm |||||$\cdot$||| | ⋅ | | in the sense that
+The policy function is $L_{z}-$ Lipschitz w.r.t some norm |||||$\cdot$||| | ⋅ | | in the sense that
 $\forall Z_{1}(\cdot|x),Z_{2}(\cdot|x)\in\mathbb{Z},\,\,|\pi(\cdot|s,Z_{1}(\cdot|x);\theta)-\pi(\cdot|s,Z_{2}(\cdot|x);\theta)|\leq L_{z}\cdot|Z_{1}(\cdot|x)-Z_{2}(\cdot|x)|$ (3)
@@
 Then, Eq.
-(4) can be interpreted as a unified training objective for both the task encoder and the policy.
-Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage iteration framework with the first stage optimizing the task encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage optimizing the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$.
+(4) can be interpreted as a unified training objective for both the context encoder and the policy.
+Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage alternating framework with the first stage updating the context encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage updating the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$.
 Hence, the performance improvement guarantee can be achieved.
-The following theorem induces a tractable upper bound for $|J^{*}(\theta)-J(\theta)|$, which forms a bedrock to connect the bridge between the intuition of prior works and the theoretical justification concerning the performance improvement guarantee.
+The following theorem induces a tractable upper bound for $|J^{*}(\theta)-J(\theta)|$, connecting the bridge between the intuition of previous works and the theoretical justification concerning the performance improvement guarantee.
 Theorem 4.3 (Return bound in COMRL).
-Denote $Z_{mutual}(\cdot|x)$ is the optimal task representation distribution served for maximizing $I(Z;M)$.
+Denote $Z_{mutual}(\cdot|x)$ as the optimal task representation distribution served for maximizing $I(Z;M)$  [^1] Whether the solution of maximizing $I(Z;M)$ is equivalent to the solution of maximizing the expected return is still an open problem.
+Hence, we introduce an extra notation here..
 Assume the reward function is upper-bounded by $R_{max}$.
@@
 Theorem 4.3 and Eq.
-(6) indicate the following two principles for the task encoder and the policy.
-For the task encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$.
-For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$ as the task representation is irrelevant for policy optimization.
+(6) indicate the following two principles for the context encoder and the policy.
+For the context encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$.
+For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$ as the task representation related terms are constants for policy optimization.
 To maximize $J(\theta)$, we can adopt standard offline RL algorithms [wu2019behavior, fujimoto2021minimalist, kumar2020conservative] similar to [yang2022unified].
@@
 As maximizing $I(Z;M)$ can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$, we argue that prior COMRL works can be interpreted by this optimization framework, thus filling the gap between the intuition and the theoretical justification.
-However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, leaving the task representation shift between two consecutive updates alone.
-In the next section, we will show that the impacts of task representation shift may lead to performance improvement collapse.
+However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, leaving the variation of the task representation alone.
+In the next section, we will show that this may lead to performance improvement collapse.
 4.2 Monotonic Performance Improvement Concerning Task Representation Shift
@@
 Definition 4.4 (Performance difference bound in COMRL).
-Denote $J^{1}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the task encoder and the policy.
-Similarly, denote $J^{2}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the task encoder and the policy.
+Denote $J^{1}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the context encoder and the policy.
+Similarly, denote $J^{2}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the context encoder and the policy.
 The performance difference bound in the COMRL setting can be defined as
 $J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq C$ (7)
-when $C$ is non-negative, the algorithm allows a monotonic performance improvement in the real environment.
+when $C$ is non-negative, the algorithm allows a monotonic performance improvement.
 According to Definition 4.4, we need to find the positive $C$ to improve the performance monotonically.
-To achieve this, we transform Eq.
-(7) to the following formulation, where $J^{mutual}(\theta_{i}),i\in[1,2]$ denote the expected return of the policy $\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{i})$ among the expectation of tasks.
-$J^{2}(\theta_{2})-J^{1}(\theta_{1})$ $=(J^{2}(\theta_{2})-J^{mutual}(\theta_{2}))+(J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1}))$
-$+(J^{mutual}(\theta_{1})-J^{2}(\theta_{1}))+(J^{2}(\theta_{1})-J^{1}(\theta_{1}))$ (8)
-Then, we introduce a mild assumption for the task representation distribution.
+We denote $J^{mutual}(\theta_{i}),i\in[1,2]$ as the expected return of the policy $\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{i})$ among the expectation of tasks and then introduce a mild assumption for the task representation distribution.
 Assumption 4.5.
@@
 $\epsilon^{mutual}_{12}\triangleq J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})>0$ (8)
-By leveraging Eq. (4.2) and the above assumptions, we derive the lower bound of the performance difference.
+By leveraging Definition 4.4 and the above assumptions, we derive the lower bound of the performance difference.
 Theorem 4.6 (Lower bound of performance difference in COMRL).
-Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of performance difference in COMRL can be formalized as:
+Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of the performance difference in COMRL can be formalized as:
 $J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (9)
 Proof: See Section 8.2 in the Appendix. ∎
-For the need of monotonic performance improvement guarantee, we further let the right-hand side of Eq. (10) be positive. Then, we would have the following requirement to achieve monotonicity.
+For the need of the monotonic performance improvement guarantee, we further let the right-hand side of Eq. (9) be positive. Then, we would have the following condition to achieve monotonicity.
 $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (10)
-The above equation unveils the key influence of the task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ for two aspects.
-If the task representation shift is too small, $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ may greater than the right-hand side of Eq.
+The above equation unveils the key influence of the variation of task representation $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ for two aspects.
+If $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ is too small, $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ may greater than the right-hand side of Eq.
 (10).
-On the contrary, if the task representation shift is too large, the right-hand side of Eq.(11) may be negative.
-Both of these cases would lead to performance improvement collapse.
-Thus, we argue that the impacts of task representation shift cannot be ignored.
-To give an example for this problem, we introduce the following assumptions.
+On the contrary, if $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ is too large, the right-hand side of Eq.(10) may be negative.
+Both of these cases would lead to the performance improvement collapse.
+Thus, we argue that the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ cannot be ignored.
+To highlight the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, we name this issue task representation shift.
+Under some mild assumptions, we can conclude that it is possible to achieve monotonic performance improvement on the condition of proper context encoder updates, as shown in Theorem 4.10.
 Assumption 4.7.
@@
 Assumption 4.9.
-The i.i.d samples from $Z_{mutual}(\cdot|x)$ can be obtained.
-Assumption 4.10.
-The task encoder is trained by fitting an empirical distribution on i.i.d samples from $Z_{mutual}(\cdot|x)$.
-Though sampling from $Z_{mutual}(\cdot|x)$ seems impractical, we can approximate it as sampling from a task encoder pre-trained by various bounds of $I(Z;M)$.
-Corollary 4.11 (Monotonic performance improvement guarantee on empirical distribution).
-Denote $\kappa$ as $\frac{2R_{max}}{(1-\gamma)^{2}}$.
-Given that the task encoder has already been trained by $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$.
-When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8, Assumption 4.9 and Assumption 4.10, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the task encoder from extra $k$ i.i.d samples, where:
-$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (12)
+There exists an $\alpha$ for any given $Z_{1}(\cdot|x)$, $|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|^{2}\leq\frac{\alpha}{b}$, where $Z_{2}(\cdot|x)$ denotes the context encoder updated by maximizing $I(Z;M)$ with the data size $b$ from the training dataset based on $Z_{1}(\cdot|x)$ and $Z^{empirical}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $b$ i.i.d samples from $Z_{mutual}(\cdot|x)$ based on $Z_{1}(\cdot|x)$  [^2] We set discrepancy to be subject to an inverse proportional function is because as the training samples increase, $Z_{2}$ and $Z^{empirical}_{2}$ should become similar..
+Theorem 4.10 (Monotonic performance improvement guarantee on training process).
+Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex.
+Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent.
+When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8 and Assumption 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder with maximizing $I(Z;M)$ from at least extra $k$ samples, where:
+$k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (11)
 Here, $\xi\in[0,1]$ is a constant.
 Proof: See Section 8.2 in the Appendix. ∎
+Theorem 4.10 shows the connection between the needed update data size $k$ and the performance improvement brought only by the policy update $\epsilon^{mutual}_{12}$.
+If the calculated $k$ is larger than the given batch size, the context encoder should avoid this update and wait for the accumulation of $\epsilon^{mutual}_{12}$.
+As $\epsilon^{mutual}_{12}$ increases, $k$ decreases.
+When $k$ is smaller than the given batch size, the monotonic performance improvement can be achieved by updating the context encoder.
+Notice that this updating process can be performed multiple times, as long as satisfying Assumption 4.7.
+The updating process is visualized in Figure 1 right.
+Compared to the original alternating framework, Theorem 4.10 unveils that the core part for better performance is to adjust the update of the context encoder to rein in the task representation shift, thereby showcasing the advantage over the previous works.
+The general algorithmic framework is shown in Algorithm 1, where the way to adjust the update of the context encoder is colored by red.
+Algorithm 1 General Algorithmic Framework Towards Reining In The Task Representation Shift
+0:  Offline training datasets $\bm{X}$, initialized policy $\pi_{\theta}$, context encoder $q_{\phi}$ and given task representation shift threshold $\beta$
+1:  for iter in alternating iterations do
+2:     // Update the context encoder
+3:     if Conditions for updating the context encoder are met then
+4:        while Accumulated task representation shift is less than $\beta$ do
+5:           Sample context from $\bm{X}$
+6:           Obtain task representations from $q_{\phi}$ with inputting the sampled context
+7:           Compute $\mathcal{L}_{\text{encoder}}$ concerning maximizing $I(Z;M)$
+8:           Update $\phi$ to minimize $\mathcal{L}_{\text{encoder}}$
+9:        end while
+10:     end if
+11:     // Update the policy
+12:     Detach the task representations
+13:     Sample training data from $\bm{X}$
+14:     Compute $\mathcal{L}_{\text{policy}}$ via standard offline RL algorithms
+15:     Update $\theta$ to minimize $\mathcal{L}_{\text{policy}}$
+16:  end for
 4.3 Practical Implementation
-To simplest make it practical, we first transform Eq. (11) to an optimization problem.
-$\min|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|,\,\,s.t.\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|>0$ (13)
-As alluded in Section 4.1, many contrastive learning-based algorithms involved in prior works [li2020focal, yuan2022robust, gao2024context, li2024towards] can be adopted to minimize $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$.
-Unfortunately, no one among them is free from parameter-tuning.
-We propose to use a classifier-based manner to address this concern.
-Specifically, after extracting from the task encoder, the task representation would be passed through a fully connected layer to get the probabilities for each task.
-Then, the task encoder is trained by the supervision of the task label and back-propagation of the cross-entropy loss.
-Note that this classifier-based manner enjoys tuning-free property and also can be seen as an approximation of $I(Z;M)$.
-To the best of our knowledge, we are the first to propose adopting a classifier-based manner to train the task encoder in the COMRL setting.
-For reining task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, since the values lying in the constraint of Eq.
-(13) are intractable, approximately estimating them may introduce a series of hyper-parameters thus significantly increasing the complexity of the algorithm.
-Instead, one insight from Corollary 4.11 is that tuning the update of the task encoder may bring better performance improvements.
-Through the experiment, we find that adjusting the update frequency of the task encoder to half of the origin is sufficient to beat the existing SOTA algorithms.
-Empirical results demonstrate that the previously ignored task representation shift does make a remarkable difference.
-Our main algorithm is shown in Algorithm 1.
-Compared to the backbone, RETRO only adds one line of code, which is colored by red.
-To maintain the consistency of prior COMRL works, we also adopt BRAC [wu2019behavior] to train the policy.
-Moreover, combining prior works with our approach to reining task representation shift is also ablated in the experiments.
-Algorithm 1 RETRO Framework
-0:  Offline Datasets $\bm{X}=\{x^{i}\}_{i=1}^{N_{train}}$, training tasks $\bm{M}=\{m^{i}\}_{i=1}^{N_{train}}$, initialized learned policy $\pi_{\theta}$, Q function $Q_{\psi}$, context encoder $q_{\phi}$, task batch size $N_{tbatch}$
-1:  while not done do
-2:     Sample task $\{m^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{M}$ and the corresponding replay buffer =$\mathcal{R}=\{x^{j}\}_{j=1}^{N_{tbatch}}\sim\bm{X}$
-3:     for step in each iteration do
-4:        Sample context $\{\bm{c}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
-5:        Obtain task representations $\{\bm{z}^{j}\sim q_{\phi}(\bm{z}|\bm{c}^{j})\}_{j=1}^{N_{tbatch}}$
-6:        // Rein the task representation shift
-7:        if step % 2 == 0 then
-8:           Compute $\mathcal{L}_{\text{RETRO}}=\textit{CrossEntropyLoss}(z^{j},m^{j})$
-9:           Update $\phi$ to minimize $\mathcal{L}_{\text{RETRO}}$
-10:        end if
-11:        Detach the task embeddings $\{\bm{z}^{j}\}_{j=1}^{N_{tbatch}}$
-12:        Sample training data $\{\bm{d}^{j}\}_{j=1}^{N_{tbatch}}\sim\mathcal{R}$
-13:        Compute $\mathcal{L}_{\text{BRAC}}$ to update $\theta,\psi$
-14:     end for
-15:  end while
+To show the impacts of the task representation shift issue, we set different manners to adjust the context encoder.
+We denote the given batch size as $bs$ and the accumulation steps of task representation shift as $acc$.
+Our manners cover 1) $k=2\times bs,acc=1$, 2) $k=3\times bs,acc=1$, 3) $k=1\times bs,acc=2$, and 4) $k=1\times bs,acc=3$.
+All manners as well as the original manner, namely $k=1\times bs,acc=1$ are performed for 8 different random seeds.
+With respect to the specific context encoder learning algorithms concerning maximizing $I(Z;M)$, we choose 3 representative algorithms that are also widely adopted in previous COMRL endeavors.
+The reconstruction-based algorithm, passing the state, action, and task representation through the decoder to reconstruct the next state and the reward signal, is applied in [zintgraf2019varibad, dorfman2021offline, li2024towards] and is proven to be the lower bound of $I(Z;M)$.
+The contrastive-based algorithm, using the task representation to compute distance metric learning loss, is applied in [li2020focal, gao2024context, li2024towards] and is proven to be the upper bound of $I(Z;M)$ as well as contain the lower bound of $I(Z;M)$.
+The cross-entropy-based algorithm is a straightforward discrete approximation to $I(Z;M)$ and is proposed by our work.
+Specifically, after extracting from the context encoder, the task representation would be passed through a fully connected layer to get the probabilities for each task.
+Then, the context encoder is trained by the supervision of the task label and back-propagation of the cross-entropy loss.
+The implementation details of these algorithms are shown in Appendix Section 8.3.
+With respect to the policy learning algorithm, to maintain the consistency of previous COMRL works, we directly adopt BRAC [wu2019behavior] to train the policy.
 5 Experiment
-Our experiments are conducted to answer 4 questions: 1) Compared to existing SOTA algorithms, can RETRO achieve more stable performance improvements and higher asymptotic performance?
-2) Can RETRO be integrated into prior works like UNICORN and FOCAL?
-3) How does reining the task representation shift with different frequencies affect RETRO?
-4) Compared to existing SOTA algorithms, can RETRO achieve lower training-time consumption?
+Our experiments are conducted to show that the previous optimization framework that ignores the impacts of task representation shift is not sufficient.
+We hope to illustrate the necessity of reining in the task representation shift, laying the foundation for further research towards better asymptotic performance of COMRL.
 5.1 Environments Settings
 We adopt MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta] benchmarks to evaluate the algorithms.
-We choose 4 competitive baselines to validate our superiority.
-The baselines cover UNICORN [li2024towards], which adopts reconstruction loss to balance the impacts of spurious correlation, CSRO [gao2024context], which adopts CLUB loss to upper-bound the impacts of the behavior policy, FOCAL [li2020focal], which directly applies distance metric learning to train the task encoder.
-and Classifier, which is our backbone without reining the task representation shift.
-The implementation details of all algorithms are shown in Appendix Section 7.3.
-Note that all these baselines can be categorized into the return discrepancy scheme stated in Section 4.1.
-Following the protocol of UNICORN, we randomly sample 20 training tasks and 20 testing tasks from the task distribution.
+Following the protocol of UNICORN  [li2024towards], we randomly sample 20 training tasks and 20 testing tasks from the task distribution.
 We train the SAC [haarnoja2018soft] agent from scratch for each task and use the collected replay buffer as the offline dataset.
 During the meta-testing phase, we adopt a fully offline setting in a few-shot manner like [li2020focal] that randomly samples a trajectory from the dataset as the context to obtain the task representation and then passes it to the meta-policy to complete the testing in the true environments.
-5.2 Baseline Comparisons
-Figure 1: Testing returns of RETRO against baselines on six benchmarks.
-Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
-Figure 1 illustrates the learning curves of RETRO against other baselines on 20 testing tasks for each benchmark.
-The results showcase that RETRO can unanimously achieve SOTA asymptotic performance across all benchmarks.
-Compared to the backbone Classifier, with adding only one line of code to rein the task representation shift, RETRO achieves remarkable performance improvements (e.g.
-approximately 20% relative performance improvements on Ant-Dir and 25% relative performance improvements on Dial-Turn).
-It is worth noting that RETRO shows comparably stable performance improvements in all the benchmarks especially on Point-Robot and Dial-Turn.
-Recall that RETRO only needs to adjust the update frequency to half of its origin, thus leading to a tuning-free algorithm.
-On the contrary, both UNICORN and CSRO are especially sensitive to the hyper-parameter settings.
-This phenomenon is particularly evident in the extremely challenging benchmark Humanoid-Dir, which holds the dimension of state space 376.
-We argue that without carefully parameter-tuning trials, UNICORN and CSRO cannot exhibit the desired performance.
-Admittedly, though more advanced algorithms may exist to better control the task representation shift, adopting the simplest manner to achieve SOTA performance makes RETRO more applicable to other fields.
-5.3 Can RETRO Be Integrated Into Other Methods?
-According to our analysis in Section 4.1, prior COMRL works ignore the impacts of task representation shift, which may lead to performance collapse.
-As our main focus is to rein the task representation shift, which is orthogonal to the prior works, RETRO should bring better performance improvements when being integrated into prior works.
-We denote RETRO-UNICORN and RETRO-FOCAL as combining the reining method into the original UNICORN and FOCAL respectively.
-To ensure fairness, we also adjust the update frequency of the task encoder in RETRO-UNICORN and RETRO-FOCAL to half of the origin.
-Figure 2: Testing returns of RETRO-UNICORN and RETRO-FOCAL against the origin on three benchmarks.
-Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
-Figure 2 demonstrates the learning curves of RETRO-UNICORN and RETRO-FOCAL against their backbones on 20 testing tasks for each benchmark.
-The results illustrate that adopting RETRO enables better performance improvements across all the benchmarks, thereby validating our core claim on reining the impacts of task representation shift.
-Note that RETRO(-Classifier) has comparable or even higher performance than RETRO-UNICORN, we believe that RETRO(-Classifier) should be a preferred option due to the tuning-free property.
-Though our goal is to design the simplest algorithm to prove the effectiveness of reining the task representation shift, we can also notice that the convergence speed is relatively lower than the backbones at the beginning of the training process.
-Hence, in the future, we envision designing a strong algorithm to control the task representation shift delicately, where the springboard is RETRO.
-5.4 The Influence Of Different Update Frequencies
-Figure 3: Testing returns of difference update frequency on two benchmarks.
-Solid curves refer to the mean performance of trials over 4 random seeds, and the shaded areas characterize the standard deviation of these trials.
-Figure 4: Average training time consumption of RETRO against other baselines on six benchmarks.
-Figure 4 illustrates the influence of different update frequencies on the performance of RETRO.
-The results show that setting the update frequency to half of the origin can generally strike a balance among the convergence speed, training stability and asymptotic performance across all the benchmarks.
-Hence, we argue that adopting RETRO as a tuning-free algorithm is acceptable.
-5.5 Can RETRO Reduce Training-Time Consumption?
-We report the training-time consumption of RETRO against other baselines across all benchmarks in Figure 4.
-All algorithms are trained with the same architecture on two RTX3090 graphic cards.
-With the exception of Humanoid-Dir, which can only run one seed per graphic card due to the limitation, the reported times for the remaining benchmarks pertain to running two seeds per card and a total of four seeds concurrently during training.
-Figure 4 shows that RETRO has a significant time-consumption reduction against UNICORN [li2024towards], CSRO [gao2024context] and FOCAL [li2020focal].
-Though this fact is based on our choice of backbone to be a classifier, it is the one proposed for the first time in our paper aiming at achieving a tuning-free algorithm.
-Compared to the backbone classifier, RETRO also achieves lower training-time consumption, as RETRO reduces the update frequency of the task encoder.
+5.2 MAIN RESULTS
+Figure 2: Testing returns of different manners to rein in the task representation shift against the original manner concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$.
+Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
+Figure 2 illustrates the learning curves of our different manners to rein in the task representation shift against the original manner on 20 testing tasks for each benchmark concerning 3 widely adopted objectives for maximizing $I(Z;M)$.
+We find that these specific objectives hold unanimous performance demonstration on all benchmarks that the previous training framework, namely $k=1\times bs,acc=1$, cannot provide the best asymptotic performance.
+These objectives focus solely on the optimization target of the task representation, namely the optimal solution of $I(Z;M)$.
+While they are guaranteed to produce reasonable results, as shown in Section 4.1, ignoring the impacts of task representation shift still hinders the capability of these algorithms to reach their full potential.
+As shown by Figure 2, even with minor changes to the algorithms, the performance improvements are substantial.
+Figure 3: Testing returns of different manners to rein in the task representation shift against the original manner concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$ on different data qualities of Ant-Dir.
+Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
+5.3 Can The Results Show Consistency Across Different Data Qualities?
+To make our claim more universal, we conduct the experiment on different data qualities.
+We collect three types of datasets with size being equal to the dataset used in Section 5.2.
+Depending on the quality of the behavior policies, we denote these datasets as random, medium, and expert respectively.
+Figure 3 demonstrates the learning curves of our different manners to rein in the task representation shift against the original manner concerning 3 widely adopted objectives for maximizing $I(Z;M)$ on Ant-Dir.
+We find that the contrastive-based algorithm fails on the random dataset.
+Except for this case, the others show consistent results that reining in the task representation shift may lead to better performance improvements.
+Based on these impressive results, we believe that developing a smarter algorithm to control the task representation shift automatically is an appealing direction and we leave this to future work.
+6 Discussion
+This section seeks to raise some interesting issues derived from our work. We hope these issues can further enable pondering on the significance of the task representation shift.
+Figure 4: Testing returns of the pre-training scheme against training from scratch of cross-entropy, contrastive, and reconstruction based objectives on Ant-Dir.
+Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
+6.1 Can The Pretraining Scheme Be Adopted To Achieve Better Performance Improvement?
+According to our analysis, the task representation shift issue only happens in the case that the context encoder needs to be trained from scratch.
+Hence, a question is raised naturally.
+If we train the context encoder in advance and use this pre-trained context encoder to train the policy directly, can we achieve better performance improvement?
+To answer this question, we conduct the experiment on Ant-Dir and still use the cross-entropy, contrastive, and reconstruction based objectives to pre-train the context encoder.
+As shown in Figure 4, there is a significant performance gap between pre-training and training from scratch.
+To better theoretically interpret this, we introduce the following Corollary.
+Corollary 6.1 (Monotonic performance improvement condition for pre-training scheme).
+Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training.
+When meeting Assumption 4.1, Assumption 4.5, the monotonic performance improvement condition for pre-training scheme is:
+$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (12)
+Proof: See Section 8.2 in the Appendix. ∎
+According to Corollary 6.1, pre-training cannot actually achieve monotonic performance improvements as $|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|$ is not equal to 0.
+As the task representation cannot be changed under the pre-training scheme, compared to training from scratch, this scheme loses some degrees of freedom to improve performance.
+Additionally, this scheme can be intuitively regarded as degrading to the problem that how to design a pre-training algorithm to better approximate $I(Z;M)$.
+Since the cross-entropy based objective is the direct approximation of $I(Z;M)$, holding better performance under the pre-training scheme is also in line with expectation.
+This analysis further enhances the importance of considering the impacts stemming from the variation of the task representation in the alternating process.
+Nevertheless, how to achieve better performance under the pre-training scheme is also an interesting problem.
+Figure 5: The 2D projection of the learned task representation space in Ant-Dir.
+Points are uniformly sampled from the evaluation datasets.
+Tasks of given goals from 0 to 6 are mapped to rainbow colors, ranging from purple to red.
+6.2 Can The Visualization Of Task Representation Be Strongly Relied Upon to Imply The Asymptotic Performance?
+Previous COMRL endeavors [li2020focal, yuan2022robust, gao2024context] mostly apply t-SNE on the learned task representation to show the differentiation of the tasks.
+They hold the belief that better performance corresponds to better differentiation of the tasks.
+However, based on our analysis, the visualization result only depends on the task representation at convergence and ignores the task representation variation during the whole optimization process.
+According to the illustration shown in Figure 5, less desirable differentiation results can also lead to better performance.
+Hence, it is insufficient to imply performance based on such evaluation principles.
+Nevertheless, we recognize that the visualization result can be seen as an auxiliary metric to assist in determining the task representation.
 7 Conclusion & Limitation
-Finding that the theoretical justification of maximizing $I(Z;M)$ can lead to performance improvement has been lacking, we constructively introduce the return discrepancy scheme into the COMRL setting, proving that maximizing $I(Z;M)$ can be interpreted as consistently raising the lower bound of the expected return for a given policy conditioning on the optimal task representation.
-However, after scrutinizing this optimization framework, we find it ignores the task representation shift between two consecutive updates, which may lead to performance improvement collapse.
-To address this problem, we turn to use the framework of performance difference bound, proving that by reining the task representation shift, it is possible to achieve monotonic performance improvements.
-To make it practical, we propose a tuning-free algorithm RETRO with only adding one line of code compared to the backbone.
-Extensive experiments have validated the superiority of RETRO on training stability, asymptotic performance and training-time consumption.
-Though we choose classifier-based loss to train the task encoder, the idea that reins the task representation shift is general, which is shown in Section 5.3.
+Finding that the performance improvement guarantee of the COMRL training framework remains underexplored, we constructively introduce the return discrepancy scheme into the COMRL setting, proving that the previous optimization framework can be linked with the general RL objective of maximizing the expected return.
+Furthermore, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation.
+Based on this finding, we propose a new issue called task representation shift and prove that it is possible to achieve monotonic performance improvements with proper context encoder updates.
+With minor changes to the previous training framework for reining in the task representation shift, the performance improvements can be substantial.
 We prospect that deeply exploring the role of task representation shift can make a profound difference in the COMRL setting.
-Naturally, one limitation is that there may exist more advanced algorithms to control the task representation shift, which is also stated in Section 5.3.
+Naturally, one limitation is that there may exist more advanced algorithms to control the task representation shift.
 […]
@@
 $|V^{\pi_{2}}_{M_{2}}-V^{\pi_{1}}_{M_{1}}|\leq 2R_{max}(\frac{\epsilon_{\pi}}{(1-\gamma)^{2}}+\frac{\gamma}{(1-\gamma)^{2}}\epsilon^{M_{2}}_{M_{1}})$ (13)
+Proof:
+$\begin{split}|V^{\pi_{2}}_{M_{2}}-V^{\pi_{1}}_{M_{1}}|&=|\sum^{\infty}\limits_{t=0}\gamma^{t}\sum\limits_{s,a}(p^{\pi_{2}}_{t,M_{2}}(s,a)-p^{\pi_{1}}_{t,M_{1}}(s,a))r(s,a)|\\ &\leq R_{max}\sum^{\infty}\limits_{t=0}\gamma^{t}\sum\limits_{s,a}|p^{\pi_{2}}_{t,M_{2}}(s,a)-p^{\pi_{1}}_{t,M_{1}}(s,a)|\\ &=2R_{max}\sum^{\infty}\limits_{t=0}\gamma^{t}D_{TV}(p^{\pi_{1}}_{t,M_{1}}(s,a)||p^{\pi_{2}}_{t,M_{2}}(s,a))\\ \end{split}$ (14)
+According to Theorem 2 (Return Bound) in [zhang2023how],
+$\sum^{\infty}\limits_{t=0}\gamma^{t}D_{TV}(p^{\pi_{1}}_{t,M_{1}}(s,a)||p^{\pi_{2}}_{t,M_{2}}(s,a))\leq(\frac{\epsilon_{\pi}}{(1-\gamma)^{2}}+\frac{\gamma}{(1-\gamma)^{2}}\epsilon^{M_{2}}_{M_{1}})$ (15)
+We bring this result back and can draw the conclusion.
 Lemma 8.2 (Inequility for $L_{1}$ deviation of the empirical distribution.[weissman2003inequalities]).
@@
 $Pr(|P-\hat{P}|\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp{(-m\epsilon^{2}/2)}$ (16)
+Proof:
+According to the inequality for $L_{1}$ deviation of the empirical distribution [weissman2003inequalities], they conclude that $Pr(|P-\hat{P}|_{1}\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp(-m\phi(\pi_{P})\epsilon^{2}/4)$, where $\phi(p)=\frac{1}{1-2p}\log\frac{1-p}{p}$ and $\pi_{P}=\max_{A\in\mathcal{A}}\min(P(A),1-P(A))$.
+The paper [weissman2003inequalities] point out that $0\leq\pi_{P}\leq 1/2,\forall P$.
+Based on this, we can derive the function $\phi(p)$ and find that $\phi(p)$ is monotonically decreasing on $(0,1/2)$.
+Therefore, we can get $\phi(p)\geq\phi(1/2)=2$.
+Then, we bring this result back and can conclude that $Pr(|P-\hat{P}|_{1}\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp(-m\epsilon^{2}/2)$.
 8.2 Missing Proofs
@@
 $+|\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]|]$ (21)
-$=\mathbb{E}_{m,x}[|\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)-\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)|$
-$+|\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)-\sum_{t=0}^{\infty}\gamma^{t}\sum_{s,a}P^{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}_{t}(s,a)R_{m}(s,a)|]$ (21)
+$=\mathbb{E}_{m,x}[|V^{\pi(\cdot|s,Z(\cdot|x);\theta)}_{m}-V^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{m}|+|V^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{m}-V^{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}_{m}|]$ (22)
 $\leq\mathbb{E}_{m,x}[\frac{2R_{max}}{(1-\gamma)^{2}}\max_{s}D_{TV}(\pi(\cdot|s,Z(\cdot|x);\theta)||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta))$
@@
 $J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (35)
-Corollary 7.5 (Monotonic performance improvement guarantee on empirical distribution).
-Denote $\kappa$ as $\frac{2R_{max}}{(1-\gamma)^{2}}$.
-Given that the task encoder has already been trained by $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$.
-When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8, Assumption 4.9 and Assumption 4.10, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the task encoder from extra $k$ i.i.d samples, where:
-$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (35)
+Theorem 8.5 (Monotonic performance improvement guarantee on training process).
+Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex.
+Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent.
+When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8 and Assumption 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder with maximizing $I(Z;M)$ from at least extra $k$ samples, where:
+$k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (36)
 Here, $\xi\in[0,1]$ is a constant.
-Proof: Let $Z_{1}$ has already been trained on $N$ i.i.d samples from $Z_{mutual}(\cdot|x)$.
-Given that to get $Z_{2}$, we need to train $Z_{1}$ by extra $k$ i.i.d samples from $Z_{mutual}(\cdot|x)$ to get the monotonic performance improvement guarantee.
-By applying Lemma 7.2, the $L_{1}$ deviation of the empirical distribution $Z_{2}(\cdot|x)$ and true $Z_{mutual}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
-$Pr(|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|<=\epsilon)\geq 1-(2^{|\mathcal{\mathbb{Z}}|}-2)\exp{(-(N+k)\epsilon^{2}/2)}$ (36)
+Proof:
+Let $Z_{1}$ have already been trained by maximizing $I(Z;M)$ to some extent.
+Given that to get $Z_{2}$, we need to train $Z_{1}$ by extra $k$ samples from the training dataset by maximizing $I(Z;M)$ to get the monotonic performance improvement guarantee.
+We begin with the following inequality:
+$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|+|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ (37)
+where $Z^{empirical}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $k$ i.i.d samples from $Z_{mutual}(\cdot|x)$ based on $Z_{1}(\cdot|x)$.
+According to Assumption 4.9, we have:
+$|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|\leq\sqrt{\frac{\alpha}{k}}$ (38)
+Now, we move our focus on solving $|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$.
+By applying Lemma 8.2, the $L_{1}$ deviation of the empirical distribution $Z^{empirical}_{2}(\cdot|x)$ and true $Z_{mutual}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
+$Pr(|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\geq\epsilon)\leq(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (39)
+$Pr(|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|<=\epsilon)\geq 1-(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (40)
 Then for a fixed $x$, with probability greater than $1-\xi$, we have:
-$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{N+k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (37)
-Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
-$k=\frac{8\kappa^{2}L_{z}^{2}}{(\epsilon^{mutual}_{12}-\kappa L_{z}\beta)^{2}}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}-N$ (38)
+$|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (41)
+Hence, we can get:
+$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}$ (42)
+Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, and recall that $\kappa$ denotes $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
+$\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ (43)
+$k\geq\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (44)
+Since we need the least number of samples that update the context encoder, we take $k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$. ∎
+Corollary 8.6 (Monotonic performance improvement condition for pre-training scheme).
+Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training.
+When meeting Assumption 4.1, Assumption 4.5, the monotonic performance improvement condition for pre-training scheme is:
+$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (45)
+Proof: Denote $J^{pretrain}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the policy.
+Similarly, denote $J^{pretrain}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the policy.
+$J^{pretrain}(\theta_{2})-J^{pretrain}(\theta_{1})$ (46)
+$=J^{pretrain}(\theta_{2})-J^{mutual}(\theta_{2})+J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})+J^{mutual}(\theta_{1})-J^{pretrain}(\theta_{1})$ (47)
+Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1, Assumption 4.1 and Assumption 4.5 into the above formulation, we can get:
+$J^{pretrain}(\theta_{2})-J^{mutual}(\theta_{2})+J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})+J^{mutual}(\theta_{1})-J^{pretrain}(\theta_{1})$ (48)
+$\geq\mathbb{E}_{m,x}[-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{2}))]+\epsilon^{mutual}_{12}$ (49)
+$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{1}))]$ (50)
+$\geq\mathbb{E}_{m,x}[-2\kappa L_{z}|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|+\epsilon^{mutual}_{12}]$ (51)
+To get the monotonic performance improvement, if we consider the condition more strictly, we need the term within the expectation of Eq. (51) to be larger than 0. Hence, we can get:
+$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{1}{2\kappa L_{z}}\epsilon^{mutual}_{12}=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (52)
 8.3 Implementation Details
-In this section, we report the implementation details of RETRO and all the baselines.
-  • FOCAL[li2020focal].
-We adopt the open-source code of FOCAL [^1] https://github.com/FOCAL-ICLR/FOCAL-ICLR/.
-To better show the performance of FOCAL in each environment, we tune the parameter $\beta$ in distance metric learning from $[0.1,0.5,1]$ and choose the best result in our baseline comparison.
-  • CSRO[gao2024context].
-We adopt the open-source code of CSRO [^2] https://github.com/MoreanP/CSRO.
-To better show the performance of CSRO in each environment, we tune the club loss weight and task representation loss weight from $[10,20,30]\times[5,10,15]$ and choose the best result in our baseline comparison.
-  • UNICORN[li2024towards].
-We obtain the source code of UNICORN from the authors.
-To better show the performance of UNICORN in each environment, we tune the parameter $\frac{\alpha\beta-1}{(1-\alpha\beta)}$ from $[0.1,0.3,0.7,1,1.5]$ and choose the best result in our baseline comparison.
-  • Classifier. Built upon the code base of FOCAL, Classifier only replaces the distance metric learning loss with cross-entropy loss, achieving tuning-free property.
-  • RETRO. Built upon the code base of Classifier, RETRO only adjusts the update frequency of the task encoder to half of the origin, by adding merely one row of code.
+In this section, we report the implementation details of all the objectives in our main paper.
+  • Contrastive-based [li2020focal]. We adopt the open-source code of FOCAL [^3] https://github.com/FOCAL-ICLR/FOCAL-ICLR/.
+  • Reconstruction-based [li2024towards]. We obtain the source code of UNICORN from the authors and adopt UNICORN-0 as the reconstruction-based objective training framework.
+  • Cross-entropy-based. Built upon the code base of FOCAL, cross-entropy-based only replaces the distance metric learning loss with cross-entropy loss.
 To maintain fairness, we make sure all benchmark-irrelevant parameters are consistent, as shown in Table 1.
@@
 task encoder network size [64, 64]
-learning rate 4e-3
+learning rate 3e-4
 Table 2: Benchmark-relevant parameters setting in the training process.
-Configurations Point-Robot Humanoid-Dir Reach Ant-Dir Button-Press Dial-Turn
-dataset size 2100 9e4 3e5 9e4 3e5 3e5
-task representation dimension 5 64 5 5 5 5
+Configurations Reach Ant-Dir Button-Press Dial-Turn
+dataset size 3e5 9e4 3e5 3e5
+task representation dimension 5 5 5 5
 For each environment, we make the benchmark-relevant parameters the same for each algorithm, as shown in Table 2.

### v3

@@
 Offline meta reinforcement learning (OMRL) has emerged as a promising approach for interaction avoidance and strong generalization performance by leveraging pre-collected data and meta-learning techniques.
-Previous context-based approaches predominantly rely on the intuition that alternating optimization between the context encoder and the policy can lead to performance improvements, as long as the context encoder follows the principle of maximizing the mutual information between the task and the task representation ($I(Z;M)$) while the policy adopts the standard offline reinforcement learning (RL) algorithms conditioning on the learned task representation.
+Previous context-based approaches predominantly rely on the intuition that alternating optimization between the context encoder and the policy can lead to performance improvements, as long as the context encoder follows the principle of maximizing the mutual information between the task variable $M$ and its latent representation $Z$ ($I(Z;M)$) while the policy adopts the standard offline reinforcement learning (RL) algorithms conditioning on the learned task representation.
 Despite promising results, the theoretical justification of performance improvements for such intuition remains underexplored.
-Inspired by the return discrepancy scheme in the model-based RL field, we find that the previous optimization framework can be linked with the general RL objective of maximizing the expected return, thereby providing a feasible explanation concerning performance improvements.
-Furthermore, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating optimization process, which may lead to performance improvement collapse.
+Inspired by the return discrepancy scheme in the model-based RL field, we find that the previous optimization framework can be linked with the general RL objective of maximizing the expected return, thereby explaining performance improvements.
+Furthermore, after scrutinizing this optimization framework, we find it ignores the variation of the task representation in the alternating optimization process, which weakens the condition necessary for monotonic performance improvements, and may therefore violate the monotonicity.
 We name this issue task representation shift and theoretically prove that the monotonic performance improvements can be guaranteed with appropriate context encoder updates.
-We set different manners to rein in the task representation shift on three widely adopted training objectives concerning maximizing $I(Z;M)$ across different data qualities.
+We use different settings to rein in the task representation shift on three widely adopted training objectives concerning maximizing $I(Z;M)$ across different data qualities.
 Empirical results show that reining in the task representation shift can indeed improve performance.
-Our work opens up a new avenue for OMRL, leading to a better understanding between the performance and the task representation.
+Our work opens up a new avenue for OMRL, leading to a better understanding between the task representation and performance improvements.
 1 Introduction
@@
 They both adopt the alternating optimization framework to train the context encoder and the policy.
-However, ours considers the variation of task representation ignored by the previous training framework.
+However, ours considers the variation of task representation ignored by the previous training framework by imposing the extra condition to determine whether needs to update the context encoder.
 Intriguingly, we formalize this training framework as an alternating two-stage optimization framework and then link it with the general RL objective of maximizing the expected return.
-To be specific, maximizing $I(Z;M)$ and adopting standard offline RL algorithms can be interpreted as consistently raising the lower bound of the expected return conditioning on the optimal task representation distribution among the expectation of tasks.
+To be specific, maximizing $I(Z;M)$ and adopting standard offline RL algorithms can be interpreted as consistently raising the lower bound of the expected return conditioning on the optimal task representation distribution among the expectation of tasks (See Section 4.1).
 This is achieved by extending the return discrepancy scheme [janner2019trust] to the COMRL framework.
@@
 More importantly, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation in the alternating process.
-Variations of the task representation that are either too large or too small can lead to performance improvement collapse.
-To highlight the characteristic, we name this issue task representation shift and theoretically prove that it is possible to achieve monotonic performance improvement guarantees with proper updates.
-To show the impacts of this issue, we set different manners to rein in the task representation shift on contrastive objective adopted in [li2020focal, gao2024context, li2024towards] and reconstruction objective adopted in [zintgraf2019varibad, dorfman2021offline, li2024towards] across different data qualities.
+This would cause the optimization framework to incorrectly conclude that the monotonic performance improvement can be guaranteed with a better approximation to the optimal task representation distribution.
+By explicitly modeling the variation of task representation, we prove that it is a critical part of monotonic performance improvement.
+Without this, the monotonicity can be violated.
+To highlight the characteristic, we name this issue task representation shift and theoretically prove that it is possible to achieve monotonic performance improvement with appropriate updates (See Section 4.2).
+To show the impacts of this issue, we use different settings to rein in the task representation shift on contrastive objective adopted in [li2020focal, gao2024context, li2024towards] and reconstruction objective adopted in [zintgraf2019varibad, dorfman2021offline, li2024towards] across different data qualities.
 These two objectives are proven to maximize the approximate bounds of $I(Z;M)$ [li2024towards].
@@
 Empirical results show that reining in the task representation shift can indeed improve performance.
-Our work opens up a new avenue for OMRL, leading to a better understanding between the performance and the task representation.
+Our work opens up a new avenue for OMRL, leading to a better understanding between the task representation and performance improvements.
 2 Related Works
@@
 Acknowledging the theoretical insights from [li2024towards], prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] mainly focus on how to design a context encoder learning algorithm to better approximate $I(Z;M)$.
-However, our work moves the focus on how to get better asymptotic performance based on a given context encoder learning algorithm.
+However, our work moves the focus on how to refine the condition of monotonic performance improvements based on a given context encoder learning algorithm.
 Center around this motivation, we identify a new issue called task representation shift, which is ignored by the previous COMRL endeavors.
@@
 Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x);\theta)$.
-Naturally, there exists an optimal task representation distribution $Z^{*}(\cdot|x)$ that meets the optimal return for any policy parameter $\theta$ is determined by $\pi(a|s,Z^{*}(\cdot|x);\theta)$.
+Naturally, there exists an optimal task representation distribution $Z^{*}(\cdot|x)$ that meets the optimal expected return for any policy parameter $\theta$ is determined by $\pi(a|s,Z^{*}(\cdot|x);\theta)$.
 Hence, the objective of COMRL is formalized as:
@@
 Previous COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to optimize the context encoder by maximizing the approximate bounds of $I(Z;M)$, as shown in Theorem 3.1.
-They hold the intuition that using such kind of manner to optimize the context encoder and adopting the standard offline RL algorithms to optimize the policy can lead to performance improvements.
+They hold the intuition that using such kind of approach to optimize the context encoder and adopting the standard offline RL algorithms to optimize the policy can lead to performance improvements.
 However, the theoretical justification of performance improvement for such intuition has been less explored.
+For later use, we denote $Z_{mutual}(\cdot|x)$ as the optimal solution for maximizing $I(Z;M)$.
+Whether $Z_{mutual}(\cdot|x)$ is equivalent to $Z^{*}(\cdot|x)$ is still an open problem.
+Hence, we introduce an extra notation here to avoid confusion.
 Theorem 3.1 ([li2024towards]).
@@
 4.1 A Performance Improvement Perspective Towards Prior Works
-Our goal is to induce an optimization objective served for COMRL to get the performance improvement guarantee.
+Our goal is to arrive at an optimization objective for COMRL to get the performance improvement guarantee.
 Motivated by the return discrepancy scheme [janner2019trust] in the model-based RL field, we can similarly build a tractable lower bound for $J^{*}(\theta)$.
@@
 where $J(\theta)$ is the expected return of the policy $\pi$ among the expectation of tasks conditioning on the learned task embedding $Z(\cdot|x)$ as $J(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]]$.
-From Definition 4.2, we can know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously.
+From Definition 4.2, we know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously.
 Then, Eq.
 (4) can be interpreted as a unified training objective for both the context encoder and the policy.
-Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage alternating framework with the first stage updating the context encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage updating the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$.
+Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage alternating framework with the first stage updating the context encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage updating the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$ as $J^{*}(\theta)\geq J(\theta)-|J^{*}(\theta)-J(\theta)|$.
 Hence, the performance improvement guarantee can be achieved.
@@
 Theorem 4.3 (Return bound in COMRL).
-Denote $Z_{mutual}(\cdot|x)$ as the optimal task representation distribution served for maximizing $I(Z;M)$  [^1] Whether the solution of maximizing $I(Z;M)$ is equivalent to the solution of maximizing the expected return is still an open problem.
-Hence, we introduce an extra notation here..
-Assume the reward function is upper-bounded by $R_{max}$.
-For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
+Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
 $|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (5)
@@
 For the context encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$.
-For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$ as the task representation related terms are constants for policy optimization.
+For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$, as the task representation related terms are constants for policy optimization.
 To maximize $J(\theta)$, we can adopt standard offline RL algorithms [wu2019behavior, fujimoto2021minimalist, kumar2020conservative] similar to [yang2022unified].
-Recall that prior COMRL works adopt alternatively training the context encoder by maximizing $I(Z;M)$ and training the policy network by conditioning on the learned task representation as well as applying the standard offline RL algorithms.
-As maximizing $I(Z;M)$ can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$, we argue that prior COMRL works can be interpreted by this optimization framework, thus filling the gap between the intuition and the theoretical justification.
-However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, leaving the variation of the task representation alone.
-In the next section, we will show that this may lead to performance improvement collapse.
+Recall that previous COMRL works adopt alternatively training the context encoder by maximizing $I(Z;M)$, which can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$ and training the policy by conditioning on the learned task representation as well as applying the standard offline RL algorithms.
+Therefore, we argue that previous COMRL works can be interpreted by this framework, thus providing an explanation concerning performance improvement guarantee.
+However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, without considering the variation of $Z(\cdot|x)$.
+In the next section, we will show that this may violate the monotonicity of the performance improvements.
 4.2 Monotonic Performance Improvement Concerning Task Representation Shift
-We attempt to build a theoretical framework that can achieve monotonic performance improvement. Hence, we turn to use the framework of performance difference bound.
+In this section, we aim to demonstrate that as the previous optimization framework in Section 4.1 doesn’t model the variation of task representation explicitly, this framework would provide an insufficient condition for monotonic performance improvement.
+We first show conditions for monotonic performance improvement of this framework.
+Corollary 4.4 (Monotonic performance improvement condition for previous COMRL works).
+When meeting Assumption 4.1, the condition for monotonic performance improvement of previous COMRL works is:
+$\epsilon^{*}_{12}\triangleq J^{*}(\theta_{2})-J^{*}(\theta_{1})\geq\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z^{*}(\cdot|x)|)$ (7)
+Proof: See Section 8.2 in the Appendix. ∎
+As shown in Corollary 4.4, the monotonic performance improvement can be guaranteed with only better approximation to $Z^{*}(\cdot|x)$.
+Next, we will introduce the performance difference bound framework to model variation of task representation.
 Definition 4.5 (Performance difference bound in COMRL).
@@
 when $C$ is non-negative, the algorithm allows a monotonic performance improvement.
-According to Definition 4.4, we need to find the positive $C$ to improve the performance monotonically.
-We denote $J^{mutual}(\theta_{i}),i\in[1,2]$ as the expected return of the policy $\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{i})$ among the expectation of tasks and then introduce a mild assumption for the task representation distribution.
-Assumption 4.5.
-The support set of $Z_{i}(\cdot|x),i\in[1,2]$ covers the support set of $Z_{mutual}(\cdot|x)$.
-Assumption 4.5 usually holds in practice as when we give $z\sim Z_{mutual}(\cdot|x)$ to the task representation distribution $Z_{i}(\cdot|x)$ parameterized by the neural network, we can always get a non-zero probability.
-Given that augmented state $s^{+}=[s,z]$, by applying standard offline RL algorithms with enough samples, the policy update would lead to performance improvement when meeting Assumption 4.5, which means
-$\epsilon^{mutual}_{12}\triangleq J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})>0$ (8)
-By leveraging Definition 4.4 and the above assumptions, we derive the lower bound of the performance difference.
+According to Definition 4.5, we need to find the positive $C$ to improve the performance monotonically. To achieve this, we can derive the lower bound of the performance difference.
 Theorem 4.6 (Lower bound of performance difference in COMRL).
-Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of the performance difference in COMRL can be formalized as:
-$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (9)
+Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1, the lower bound of the performance difference in COMRL can be formalized as:
+$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (9)
 Proof: See Section 8.2 in the Appendix. ∎
 For the need of the monotonic performance improvement guarantee, we further let the right-hand side of Eq. (32) be positive. Then, we would have the following condition to achieve monotonicity.
-$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (10)
-The above equation unveils the key influence of the variation of task representation $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ for two aspects.
-If $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ is too small, $|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ may greater than the right-hand side of Eq.
-(10).
-On the contrary, if $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ is too large, the right-hand side of Eq.(10) may be negative.
-Both of these cases would lead to the performance improvement collapse.
-Thus, we argue that the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ cannot be ignored.
+$\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]\geq 0$ (10)
+Compared to the condition in Corollary 4.4, Eq.
+10 presents an additional term $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ to achieve monotonic performance improvement, which corresponds exactly to the previous ignored impacts stemming from the variation of the task representation.
+To achieve monotonic performance improvements, the optimization of task representation should consider not only the approximation to $Z^{*}(\cdot|x)$, but also the magnitude of the update.
+For example, if we assume that $Z_{1}(\cdot|x)$ is trained from scratch and the update process from $Z_{1}(\cdot|x)$ to $Z_{2}(\cdot|x)$ brings $Z_{2}(\cdot|x)$ close to $Z^{*}(\cdot|x)$, then with the condition in Corollary 4.4, the monotonic performance improvement can be easily achieved with small $\epsilon^{*}_{12}$.
+However, for the condition in Eq.
+(10), small $\epsilon^{*}_{12}$ may cause the violation of monotonicity as $|Z_{1}(\cdot|x)-Z_{2}(\cdot|x)|$ remains large.
 To highlight the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, we name this issue Task Representation Shift.
-Under some mild assumptions, we can conclude that it is possible to achieve monotonic performance improvement on the condition of proper context encoder updates, as shown in Theorem 4.10.
+Under some mild assumptions, we can conclude that it is possible to achieve monotonic performance improvement with sufficient policy improvement $\epsilon^{*}_{12}$ on the condition of appropriate encoder updates, as shown in Theorem 4.10.
 Assumption 4.7.
-The impacts of task representation shift can be upper-bounded by $\beta$.
+The impacts of task representation shift are upper-bounded by $\beta$ and less than the policy improvement $\epsilon^{*}_{12}$ with a certain coefficient.
 Assumption 4.8.
@@
 Assumption 4.9.
-There exists an $\alpha$ for any given $Z_{1}(\cdot|x)$, $|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|^{2}\leq\frac{\alpha}{b}$, where $Z_{2}(\cdot|x)$ denotes the context encoder updated by maximizing $I(Z;M)$ with the data size $b$ from the training dataset based on $Z_{1}(\cdot|x)$ and $Z^{empirical}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $b$ i.i.d samples from $Z_{mutual}(\cdot|x)$ based on $Z_{1}(\cdot|x)$  [^2] We set discrepancy to be subject to an inverse proportional function is because as the training samples increase, $Z_{2}$ and $Z^{empirical}_{2}$ should become similar..
+There exists an $\alpha$ for any given $Z_{1}(\cdot|x)$, $|Z_{2}(\cdot|x)-\tilde{Z}_{2}(\cdot|x)|^{2}\leq\frac{\alpha}{b}$, where $Z_{2}(\cdot|x)$ denotes the context encoder updated by maximizing $I(Z;M)$ with the data size $b$ randomly sampled from the training dataset based on $Z_{1}(\cdot|x)$ and $\tilde{Z}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $b$ i.i.d samples from $Z^{*}(\cdot|x)$ based on $Z_{1}(\cdot|x)$  [^1] We set discrepancy to be subject to an inverse proportional function is because as the training samples increase, $Z_{2}$ and $\tilde{Z}_{2}$ should become similar..
 Theorem 4.10 (Monotonic performance improvement guarantee on training process).
-Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex.
+Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex.
 Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent.
-When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8 and Assumption 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder with maximizing $I(Z;M)$ from at least extra $k$ samples, where:
+When meeting Assumption 4.1, 4.7, 4.8 and 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder via maximizing $I(Z;M)$ from at least extra $k$ samples, where:
 $k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (11)
@@
 Proof: See Section 8.2 in the Appendix. ∎
-Theorem 4.10 shows the connection between the needed update data size $k$ and the performance improvement brought only by the policy update $\epsilon^{mutual}_{12}$.
-If the calculated $k$ is larger than the given batch size, the context encoder should avoid this update and wait for the accumulation of $\epsilon^{mutual}_{12}$.
-As $\epsilon^{mutual}_{12}$ increases, $k$ decreases.
+Theorem 4.10 shows the connection between the needed update data size $k$ and the performance improvement brought only by the policy update $\epsilon^{*}_{12}$.
+If the calculated $k$ is larger than the given batch size, the context encoder should avoid this update and wait for the accumulation of $\epsilon^{*}_{12}$.
+As $\epsilon^{*}_{12}$ increases, $k$ decreases.
 When $k$ is smaller than the given batch size, the monotonic performance improvement can be achieved by updating the context encoder.
@@
 The updating process is visualized in Figure 1 right.
-Compared to the original alternating framework, Theorem 4.10 unveils that the core part for better performance is to adjust the update of the context encoder to rein in the task representation shift, thereby showcasing the advantage over the previous works.
+Compared to the original alternating framework, Theorem 4.10 unveils that the core part for better performance improvement is to adjust the update of the context encoder to rein in the task representation shift, thereby showcasing the advantage over the previous works.
 The general algorithmic framework is shown in Algorithm 1, where the way to adjust the update of the context encoder is colored by red.
@@
 4.3 Practical Implementation
-To show the impacts of the task representation shift issue, we set different manners to adjust the context encoder.
+To show the impacts of the task representation shift issue, we use different settings to adjust the context encoder.
 We denote the given batch size as $bs$ and the accumulation steps of task representation shift as $acc$.
-Our manners cover 1) $k=2\times bs,acc=1$, 2) $k=3\times bs,acc=1$, 3) $k=1\times bs,acc=2$, and 4) $k=1\times bs,acc=3$.
-All manners as well as the original manner, namely $k=1\times bs,acc=1$ are performed for 8 different random seeds.
+Our settings include 1) $k=2\times bs,acc=1$, 2) $k=3\times bs,acc=1$, 3) $k=1\times bs,acc=2$, and 4) $k=1\times bs,acc=3$.
+All settings as well as the original setting, namely $k=1\times bs,acc=1$ are performed for 8 different random seeds.
 With respect to the specific context encoder learning algorithms concerning maximizing $I(Z;M)$, we choose 3 representative algorithms that are also widely adopted in previous COMRL endeavors.
@@
 Our experiments are conducted to show that the previous optimization framework that ignores the impacts of task representation shift is not sufficient.
-We hope to illustrate the necessity of reining in the task representation shift, laying the foundation for further research towards better asymptotic performance of COMRL.
+We hope to illustrate the potential of reining in the task representation shift, laying the foundation for further research towards better performance improvement of COMRL.
 5.1 Environments Settings
@@
 5.2 MAIN RESULTS
-Figure 2: Testing returns of different manners to rein in the task representation shift against the original manner concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$.
+Figure 2: Testing returns of different settings to rein in the task representation shift against the original setting concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$.
 Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
-Figure 2 illustrates the learning curves of our different manners to rein in the task representation shift against the original manner on 20 testing tasks for each benchmark concerning 3 widely adopted objectives for maximizing $I(Z;M)$.
+Figure 2 illustrates the learning curves of our different settings to rein in the task representation shift against the original setting on 20 testing tasks for each benchmark concerning 3 widely adopted objectives for maximizing $I(Z;M)$.
 We find that these specific objectives hold unanimous performance demonstration on all benchmarks that the previous training framework, namely $k=1\times bs,acc=1$, cannot provide the best asymptotic performance.
@@
 Depending on the quality of the behavior policies, we denote these datasets as random, medium, and expert respectively.
-Figure 3 demonstrates the learning curves of our different manners to rein in the task representation shift against the original manner concerning 3 widely adopted objectives for maximizing $I(Z;M)$ on Ant-Dir.
+Figure 3 demonstrates the learning curves of our different settings to rein in the task representation shift against the original setting concerning 3 widely adopted objectives for maximizing $I(Z;M)$ on Ant-Dir.
 We find that the contrastive-based algorithm fails on the random dataset.
@@
 Corollary 6.1 (Monotonic performance improvement condition for pre-training scheme).
-Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training.
-When meeting Assumption 4.1, Assumption 4.5, the monotonic performance improvement condition for pre-training scheme is:
-$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (12)
+Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, the monotonic performance improvement condition for pre-training scheme is:
+$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (12)
 Proof: See Section 8.2 in the Appendix. ∎
-According to Corollary 6.1, pre-training cannot actually achieve monotonic performance improvements as $|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|$ is not equal to 0.
+Due to the implementation error and approximation error, $|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|$ can not equal to 0.
+Hence, according to Corollary 6.1, pre-training cannot actually achieve monotonic performance improvements.
 As the task representation cannot be changed under the pre-training scheme, compared to training from scratch, this scheme loses some degrees of freedom to improve performance.
-Additionally, this scheme can be intuitively regarded as degrading to the problem that how to design a pre-training algorithm to better approximate $I(Z;M)$.
+Additionally, this scheme can be intuitively regarded as degrading to the problem that how to design a pre-training algorithm to better approximate $Z^{*}(\cdot|x)$.
+Notice that there exists the decomposition $|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|\leq|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$.
 Since the cross-entropy based objective is the direct approximation of $I(Z;M)$, holding better performance under the pre-training scheme is also in line with expectation.
-This analysis further enhances the importance of considering the impacts stemming from the variation of the task representation in the alternating process.
+This analysis further enhances the importance of considering the impacts of task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ in the alternating process.
 Nevertheless, how to achieve better performance under the pre-training scheme is also an interesting problem.
@@
 7 Conclusion & Limitation
-Finding that the performance improvement guarantee of the COMRL training framework remains underexplored, we constructively introduce the return discrepancy scheme into the COMRL setting, proving that the previous optimization framework can be linked with the general RL objective of maximizing the expected return.
-Furthermore, after scrutinizing this optimization framework, we find it ignores the impacts stemming from the variation of the task representation.
-Based on this finding, we propose a new issue called task representation shift and prove that it is possible to achieve monotonic performance improvements with proper context encoder updates.
+Finding that the performance improvement guarantee of the COMRL training framework remains under-explored, we constructively provide a theoretical framework to link the training scheme in COMRL with the general RL objective of maximizing the expected return.
+After scrutinizing this framework, we further find it ignores the variation of the task representation, which weakens the condition necessary for monotonic performance improvements.
+Based on this finding, we propose a new issue called task representation shift, refine the condition for monotonicity, and prove that monotonic performance improvements can be achieved with appropriate context encoder updates.
 With minor changes to the previous training framework for reining in the task representation shift, the performance improvements can be substantial.
@@
 Naturally, one limitation is that there may exist more advanced algorithms to control the task representation shift.
+Additionally, our work focuses on the representation part for monotonic performance improvement, leaving the policy learning part alone.
+Hence, one direction that merits further research is to design a policy learning algorithm to achieve better policy improvements conditioning on the optimal task representation distribution.
+We leave these interesting problems to future work.
 […]
@@
 Theorem 8.3 (Return bound in COMRL).
-Denote $Z_{mutual}(\cdot|x)$ is the optimal task representation distribution served for maximizing $I(Z;M)$.
-Assume the reward function is upper-bounded by $R_{max}$.
-For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
+Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
 $|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (17)
@@
 Eq. (23) is the result of directly applying Lemma 8.1 and Eq. (24) is the result of directly applying Assumption 4.1. ∎
+Corollary 8.4 (Monotonic performance improvement condition for previous COMRL works).
+When meeting Assumption 4.1, the condition for monotonic performance improvement of previous COMRL works is:
+$\epsilon^{*}_{12}\triangleq J^{*}(\theta_{2})-J^{*}(\theta_{1})\geq\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z^{*}(\cdot|x)|)$ (25)
+Proof:
+$J(\theta_{2})-J(\theta_{1})=J(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J(\theta_{1})$ (26)
+Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1 and Assumption 4.1 into the above formulation, we can get:
+$J(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J(\theta_{1})$ (27)
+$\geq\mathbb{E}_{m,x}[-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z(\cdot|x);\theta_{2})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{2}))]+\epsilon^{*}_{12}$ (28)
+$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z(\cdot|x);\theta_{1})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{1}))]$ (29)
+$\geq\mathbb{E}_{m,x}[-2\kappa L_{z}|Z(\cdot|x)-Z^{*}(\cdot|x)|+\epsilon^{*}_{12}]$ (30)
+To get the monotonic performance improvement, we need Eq. (30) to be larger than 0. Hence, we can get:
+$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (31)
 Theorem 8.5 (Lower bound of performance difference in COMRL).
-Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1 and Assumption 4.5, the lower bound of performance difference in COMRL can be formalized as:
-$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (25)
-Proof:
+Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1, the lower bound of the performance difference in COMRL can be formalized as:
+$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (32)
+Proof: We can introduce the following decomposition
 $J^{2}(\theta_{2})-J^{1}(\theta_{1})$ (33)
@@
 $=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (37)
-$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z_{mutual}}+G^{\theta_{2}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{mutual}}+G^{\theta_{1}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (31)
-Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1, Assumption 4.1 and Assumption 4.5 into the above formulation, we can get:
-$G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z_{mutual}}+G^{\theta_{2}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{mutual}}+G^{\theta_{1}}_{Z_{mutual}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (32)
-$\geq-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{2}))+\epsilon^{mutual}_{12}$
-$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1}))$
+$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z^{*}}+G^{\theta_{2}}_{Z^{*}}-G^{\theta_{1}}_{Z^{*}}+G^{\theta_{1}}_{Z^{*}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (38)
+Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1 and Assumption 4.1 into the above formulation, we can get:
+$G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z^{*}}+G^{\theta_{2}}_{Z^{*}}-G^{\theta_{1}}_{Z^{*}}+G^{\theta_{1}}_{Z^{*}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (39)
+$\geq-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{2}))+\epsilon^{*}_{12}$
+$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1}))$
 $-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1}))$ (40)
-$\geq-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+\epsilon^{mutual}_{12}$
-$-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (34)
+$\geq-\kappa L_{z}|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+\epsilon^{*}_{12}$
+$-\kappa L_{z}|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (41)
 Simplifying Eq. (41) we can get:
-$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{mutual}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (35)
+$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (42)
 Theorem 8.6 (Monotonic performance improvement guarantee on training process).
-Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex.
+Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex.
 Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent.
-When meeting Assumption 4.1, Assumption 4.5, Assumption 4.7, Assumption 4.8 and Assumption 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder with maximizing $I(Z;M)$ from at least extra $k$ samples, where:
+When meeting Assumption 4.1, 4.7, 4.8 and 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder via maximizing $I(Z;M)$ from at least extra $k$ samples, where:
 $k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (43)
@@
 We begin with the following inequality:
-$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|+|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$ (37)
-where $Z^{empirical}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $k$ i.i.d samples from $Z_{mutual}(\cdot|x)$ based on $Z_{1}(\cdot|x)$.
+$|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|\leq|Z_{2}(\cdot|x)-\tilde{Z}_{2}(\cdot|x)|+|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|$ (44)
+where $\tilde{Z}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $k$ i.i.d samples from $Z^{*}(\cdot|x)$ based on $Z_{1}(\cdot|x)$.
 According to Assumption 4.9, we have:
-$|Z_{2}(\cdot|x)-Z^{empirical}_{2}(\cdot|x)|\leq\sqrt{\frac{\alpha}{k}}$ (38)
-Now, we move our focus on solving $|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|$.
-By applying Lemma 8.2, the $L_{1}$ deviation of the empirical distribution $Z^{empirical}_{2}(\cdot|x)$ and true $Z_{mutual}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
-$Pr(|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\geq\epsilon)\leq(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (39)
-$Pr(|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|<=\epsilon)\geq 1-(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (40)
+$|Z_{2}(\cdot|x)-\tilde{Z}_{2}(\cdot|x)|\leq\sqrt{\frac{\alpha}{k}}$ (45)
+Now, we move our focus on solving $|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|$.
+By applying Lemma 8.2, the $L_{1}$ deviation of the empirical distribution $\tilde{Z}_{2}(\cdot|x)$ and true $Z^{*}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
+$Pr(|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|\geq\epsilon)\leq(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (46)
+$Pr(|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|<=\epsilon)\geq 1-(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (47)
 Then for a fixed $x$, with probability greater than $1-\xi$, we have:
-$|Z^{empirical}_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (41)
+$|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (48)
 Hence, we can get:
-$|Z_{2}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}$ (42)
-Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, and recall that $\kappa$ denotes $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
-$\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}-\frac{1}{2}\beta$ (43)
+$|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}$ (49)
+Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$, and recall that $\kappa$ denotes $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
+$\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ (50)
 $k\geq\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (51)
@@
 Corollary 8.7 (Monotonic performance improvement condition for pre-training scheme).
-Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training.
-When meeting Assumption 4.1, Assumption 4.5, the monotonic performance improvement condition for pre-training scheme is:
-$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (45)
+Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, the monotonic performance improvement condition for pre-training scheme is:
+$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (52)
 Proof: Denote $J^{pretrain}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the policy.
@@
 $J^{pretrain}(\theta_{2})-J^{pretrain}(\theta_{1})$ (53)
-$=J^{pretrain}(\theta_{2})-J^{mutual}(\theta_{2})+J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})+J^{mutual}(\theta_{1})-J^{pretrain}(\theta_{1})$ (47)
-Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1, Assumption 4.1 and Assumption 4.5 into the above formulation, we can get:
-$J^{pretrain}(\theta_{2})-J^{mutual}(\theta_{2})+J^{mutual}(\theta_{2})-J^{mutual}(\theta_{1})+J^{mutual}(\theta_{1})-J^{pretrain}(\theta_{1})$ (48)
-$\geq\mathbb{E}_{m,x}[-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{2}))]+\epsilon^{mutual}_{12}$ (49)
-$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta_{1}))]$ (50)
-$\geq\mathbb{E}_{m,x}[-2\kappa L_{z}|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|+\epsilon^{mutual}_{12}]$ (51)
-To get the monotonic performance improvement, if we consider the condition more strictly, we need the term within the expectation of Eq. (51) to be larger than 0. Hence, we can get:
-$|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|\leq\frac{1}{2\kappa L_{z}}\epsilon^{mutual}_{12}=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{mutual}_{12}$ (52)
+$=J^{pretrain}(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J^{pretrain}(\theta_{1})$ (54)
+Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1 and Assumption 4.1 into the above formulation, we can get:
+$J^{pretrain}(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J^{pretrain}(\theta_{1})$ (55)
+$\geq\mathbb{E}_{m,x}[-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{2}))]+\epsilon^{*}_{12}$ (56)
+$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{1}))]$ (57)
+$\geq\mathbb{E}_{m,x}[-2\kappa L_{z}|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|+\epsilon^{*}_{12}]$ (58)
+To get the monotonic performance improvement, we need Eq. (58) to be larger than 0. Hence, we can get:
+$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (59)
 8.3 Implementation Details

### v4

@@
 Inspired by the return discrepancy scheme in the model-based RL field, we find that the previous optimization framework can be linked with the general RL objective of maximizing the expected return, thereby explaining performance improvements.
-Furthermore, after scrutinizing this optimization framework, we find it ignores the variation of the task representation in the alternating optimization process, which weakens the condition necessary for monotonic performance improvements, and may therefore violate the monotonicity.
+Furthermore, after scrutinizing this optimization framework, we observe that the condition for monotonic performance improvements does not consider the variation of the task representation.
+When these variations are considered, the previously established condition may no longer be sufficient to ensure monotonicity, thereby impairing the optimization process.
 We name this issue task representation shift and theoretically prove that the monotonic performance improvements can be guaranteed with appropriate context encoder updates.
@@
 Our work opens up a new avenue for OMRL, leading to a better understanding between the task representation and performance improvements.
+[^1] Code: https://github.com/betray12138/Task-Representation-Shift
 1 Introduction
-RL has driven impressive advances in many complex decision-making problems in recent years [silver2018general, schrittwieser2020mastering, zhang2023safe], primarily through online RL methods.
+RL has driven impressive advances in many complex decision-making problems in recent years (Silver et al., 2018; Schrittwieser et al., 2020; Zhang et al., 2023b; 2024), primarily through online RL methods.
 However, the extensive interactions required by online RL entail high costs and safety concerns, posing significant challenges for real-world applications.
-Offline RL [wu2019behavior, levine2020offline] offers an appealing alternative by efficiently leveraging pre-collected data for policy learning, thereby circumventing the need for online interaction with the environment.
-This advantage extends the application of RL, covering healthcare [fatemi2022semi, tang2022leveraging], robotics [sinha2022s4rl, pmlr-v164-kumar22a] and games [schrittwieser2021online, mathieu2023alphastar].
-Though demonstrating its superiority, offline RL holds a notable drawback towards generalizing to the unknown [ghosh2021generalization].
-As a remedy, by combining the meta-learning techniques, OMRL [li2020focal, mitchell2021offline, xu2022prompting] has emerged as an effective training scheme toward strong generalization performance and fast adaptation capability while maintaining the merits of offline RL.
-Among the OMRL research, context-based OMRL (COMRL) algorithms [li2020focal, li2024towards] hold a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs).
-Specifically, these methods [li2020focal, gao2024context, li2024towards] propose to train a context encoder via maximizing $I(Z;M)$ to learn the task representation from the collected context (e.g.
+Offline RL (Wu et al., 2019; Levine et al., 2020) offers an appealing alternative by efficiently leveraging pre-collected data for policy learning, thereby circumventing the need for online interaction with the environment.
+This advantage extends the application of RL, covering healthcare (Fatemi et al., 2022; Tang et al., 2022), robotics (Sinha et al., 2022; Kumar et al., 2022) and games (Schrittwieser et al., 2021).
+Though demonstrating its superiority, offline RL holds a notable drawback towards generalizing to the unknown (Ghosh et al., 2021).
+As a remedy, by combining the meta-learning techniques, OMRL (Li et al., 2020; Mitchell et al., 2021; Xu et al., 2022) has emerged as an effective training scheme toward strong generalization performance and fast adaptation capability while maintaining the merits of offline RL.
+Among the OMRL research, context-based OMRL (COMRL) algorithms (Li et al., 2020; 2024) hold a popular paradigm that seeks optimal meta-policy conditioning on the context of Markov Decision Processes (MDPs).
+Specifically, these methods (Li et al., 2020; Gao et al., 2024; Li et al., 2024) propose to train a context encoder via maximizing $I(Z;M)$ to learn the task representation from the collected context (e.g.
 trajectories) and train the downstream policy with standard offline RL algorithms conditioning on the task representation, as shown in Figure 1 left.
@@
 They both adopt the alternating optimization framework to train the context encoder and the policy.
-However, ours considers the variation of task representation ignored by the previous training framework by imposing the extra condition to determine whether needs to update the context encoder.
+However, our training framework considers the previously ignored variation of task representation by introducing an extra condition to decide whether the context encoder should be updated.
 Intriguingly, we formalize this training framework as an alternating two-stage optimization framework and then link it with the general RL objective of maximizing the expected return.
-To be specific, maximizing $I(Z;M)$ and adopting standard offline RL algorithms can be interpreted as consistently raising the lower bound of the expected return conditioning on the optimal task representation distribution among the expectation of tasks (See Section 4.1).
-This is achieved by extending the return discrepancy scheme [janner2019trust] to the COMRL framework.
+To be specific, maximizing $I(Z;M)$ and adopting standard offline RL algorithms can be interpreted as consistently raising the lower bound of the expected return conditioning on the optimal task representation distribution (See Section 4.1).
+This is achieved by extending the return discrepancy scheme (Janner et al., 2019) to the COMRL framework.
 Thus, this analysis provides a feasible explanation for the performance improvement guarantee.
@@
 This would cause the optimization framework to incorrectly conclude that the monotonic performance improvement can be guaranteed with a better approximation to the optimal task representation distribution.
-By explicitly modeling the variation of task representation, we prove that it is a critical part of monotonic performance improvement.
-Without this, the monotonicity can be violated.
-To highlight the characteristic, we name this issue task representation shift and theoretically prove that it is possible to achieve monotonic performance improvement with appropriate updates (See Section 4.2).
-To show the impacts of this issue, we use different settings to rein in the task representation shift on contrastive objective adopted in [li2020focal, gao2024context, li2024towards] and reconstruction objective adopted in [zintgraf2019varibad, dorfman2021offline, li2024towards] across different data qualities.
-These two objectives are proven to maximize the approximate bounds of $I(Z;M)$ [li2024towards].
-We also introduce the cross-entropy objective, which is a straightforward discrete approximation to $I(Z;M)$.
+By explicitly modeling the variation of task representation, we conclude that it is a critical part of monotonic performance improvement.
+To highlight the characteristic, we name this issue task representation shift and theoretically prove that it is possible to achieve monotonic performance improvement with appropriate context encoder updates (See Section 4.2).
+To show the impacts of this issue, we use different settings to rein in the task representation shift on three widely adopted objectives concerning maximizing $I(Z;M)$, covering the upper bound, the lower bound, and the direct approximation.
 Empirical results show that reining in the task representation shift can indeed improve performance.
@@
 2 Related Works
-Offline RL.
-Offline RL [fujimoto2019off] leverages a fully offline dataset to perform policy optimization and value function learning, avoiding explicit interaction with the environment.
-To overcome the notorious distribution shift and the induced value overestimation issues, prior works adopt methods based on policy regularization [fujimoto2019off, kumar2019stabilizing, fujimoto2021minimalist], value penalty [wu2019behavior, kumar2020conservative] and sequence modeling [chen2021decision, janner2021offline].
-While these methods have demonstrated promising performance on a single task, their generalization performance to the unknown remains limited [ghosh2021generalization].
-Our work follows the offline RL setting, attempting to adopt meta-learning techniques to improve the generalization performance.
-Meta RL.
-Meta RL has emerged as an enticing avenue for effectively adapting to unknown environments [finn2017model].
-Unlike traditional gradient-based meta RL methods [finn2017model, mitchell2021offline], context-based meta RL methods demonstrate superior advantages in both sample efficiency and asymptotic performance [rakelly2019efficient, zintgraf2019varibad, li2024towards].
-Our work delves into this promising domain to advance the capabilities of context-based meta RL further.
 Context-Based Offline Meta RL.
 As the marriage between context-based meta RL and offline RL, COMRL combines the merits of both sides.
-Specifically, COMRL methods [li2020focal, yuan2022robust, gao2024context, li2024towards] leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input.
+Specifically, COMRL methods (Li et al., 2020; Yuan & Lu, 2022; Gao et al., 2024; Li et al., 2024) leverage the offline dataset to train a context encoder to learn robust task representations, and then pass the representation to the policy and value function as input.
 At test time, COMRL methods leverage the generalization ability of the learned context encoder to perform meta-adaptation.
-Acknowledging the theoretical insights from [li2024towards], prior COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] mainly focus on how to design a context encoder learning algorithm to better approximate $I(Z;M)$.
-However, our work moves the focus on how to refine the condition of monotonic performance improvements based on a given context encoder learning algorithm.
-Center around this motivation, we identify a new issue called task representation shift, which is ignored by the previous COMRL endeavors.
+According to the theoretical insights from (Li et al., 2024), prior COMRL works (Li et al., 2020; Yuan & Lu, 2022; Gao et al., 2024; Li et al., 2024) mainly focus on designing a context encoder learning algorithm to better approximate $I(Z;M)$.
+However, our work turns the focus to refining the condition of monotonic performance improvements based on a given context encoder learning algorithm.
+Centering around this motivation, we identify a new issue called task representation shift, which is ignored by the previous COMRL endeavors.
 Performance Improvement Guarantee.
 Ensuring performance improvement is a key concern in both online and offline reinforcement learning settings.
-In online RL, performance improvement guarantees are often established through methods such as performance difference bounds [schulman2015trust, kakade2002approximately, ji2022update, zhang2023how], return discrepancy [janner2019trust, luo2018algorithmic], and regret bounds [osband2014model, curi2020efficient].
-In offline RL, CQL-based methods [kumar2020conservative, yu2021combo] also enjoy safe policy improvement guarantees.
-However, most works focus on the single-task setting, leaving the performance improvement guarantees in the COMRL settings largely unexplored.
-While ContrBAR [choshen2023contrabar] also benefits from the performance improvement guarantee in the OMRL setting, it is specifically served for its own framework.
-In contrast, our work addresses a broader class of algorithms that maximize $I(Z;M)$, including but not limited to ContrBAR.
+In online RL, performance improvement guarantees are often established through methods such as performance difference bounds (Kakade & Langford, 2002; Schulman et al., 2015; Ji et al., 2022; Zhang et al., 2023a), return discrepancy (Luo et al., 2018; Janner et al., 2019), and regret bounds (Osband & Van Roy, 2014; Curi et al., 2020).
+In offline RL, CQL-based methods (Kumar et al., 2020; Yu et al., 2021) also enjoy safe policy improvement guarantees.
+However, most works focus on the single-task setting, leaving the performance improvement guarantees in the context-based meta-RL settings largely unexplored.
+While ContrBAR (Choshen & Tamar, 2023) also benefits from the performance improvement guarantee, the theoretical insight focuses on the online setting.
+Additionally, it is tailored to one particular approximation of $I(Z;M)$ as it makes assumptions specific to this approximation.
+In contrast, our work focuses on the general offline setting, addressing a broader class of algorithms that maximize various bounds of $I(Z;M)$.
 3 Preliminaries
 3.1 Problem Statement
-A task in RL is generally formalized as a fully observed MDP [puterman2014markov], which is represented by a tuple $(\mathcal{S},\mathcal{A},P,\rho_{0},R,\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(s^{\prime}|s,a)$, reward function $R(s,a)$, initial state distribution $\rho_{0}(s)$ and discount factor $\gamma\in[0,1]$.
+A task in RL is generally formalized as a fully observed MDP (Puterman, 2014), which is represented by a tuple $(\mathcal{S},\mathcal{A},P,\rho_{0},R,\gamma)$ with state space $\mathcal{S}$, action space $\mathcal{A}$, transition function $P(s^{\prime}|s,a)$, reward function $R(s,a)$, initial state distribution $\rho_{0}(s)$ and discount factor $\gamma\in[0,1]$.
 For arbitrary policy $\pi(a|s)$, we denote the state-action distribution at timestep $t$ as $d_{\pi,t}(s,a)\triangleq\text{Pr}(s_{t}=s,a_{t}=a|s_{0}\sim\rho_{0},a_{t}\sim\pi,s_{t+1}\sim P,\forall t\geq 0)$.
@@
 The ultimate goal is to find an optimal policy $\pi(a|s)$ to maximize the expected return $\mathbb{E}_{(s,a)\sim d_{\pi}}[R(s,a)]$.
-Given a distribution of tasks $p(M)$, every task $m$ sampled from $p(M)$ is an MDP sharing the same state space and action space but differing from transition dynamics and reward functions.
-It is widely assumed that the tasks in the meta RL setting are randomly sampled from $p(M)$ [li2020focal, yuan2022robust, gao2024context, li2024towards].
+It is widely assumed that the task $m$ in the meta RL setting is randomly sampled from the task distribution $p(M)$ (Li et al., 2020; Yuan & Lu, 2022; Gao et al., 2024; Li et al., 2024).
 We denote the number of training tasks as $N$.
@@
 3.2 Context-based OMRL
-Previous COMRL endeavors hold that the variation over different tasks can be represented by a compact representation $z$ [yuan2022robust].
-In practice, they choose to train a context encoder $Z(\cdot|x)$ to extract the task information [li2020focal, dorfman2021offline], where $x$ denotes the given context.
-Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x);\theta)$.
-Naturally, there exists an optimal task representation distribution $Z^{*}(\cdot|x)$ that meets the optimal expected return for any policy parameter $\theta$ is determined by $\pi(a|s,Z^{*}(\cdot|x);\theta)$.
+Previous COMRL endeavors employ a compact representation $z$ to capture/quantify the variation over tasks (Yuan & Lu, 2022).
+In practice, they choose to train a context encoder $Z(\cdot|x;\phi)$ to extract the task information (Li et al., 2020; Dorfman et al., 2021), where $x$ denotes the given context and $\phi$ denotes the parameters of the context encoder.
+Then, the policy is learned by conditioning on the task representation as $\pi(a|s,Z(\cdot|x;\phi);\theta)$.
+We assume there exists an optimal task representation distribution $Z(\cdot|x;\phi^{*})$ that meets the optimal expected return for any policy parameter $\theta$ is determined by $\pi(a|s,Z(\cdot|x;\phi^{*});\theta)$.
 Hence, the objective of COMRL is formalized as:
-$\max_{\theta}J^{*}(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]$ (1)
+$\max_{\theta}J^{*}(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x;\phi^{*});\theta)}}[R_{m}(s,a)]]$ (1)
 where $R_{m}(s,a)$ denotes the ground-truth reward function of the task $m$.
-Previous COMRL works [li2020focal, yuan2022robust, gao2024context, li2024towards] are proven to optimize the context encoder by maximizing the approximate bounds of $I(Z;M)$, as shown in Theorem 3.1.
+Previous COMRL works (Li et al., 2020; 2021; Yuan & Lu, 2022; Gao et al., 2024; Li et al., 2024) are proven to optimize the context encoder by maximizing the approximate bounds of $I(Z;M)$, as shown in Theorem 3.1.
 They hold the intuition that using such kind of approach to optimize the context encoder and adopting the standard offline RL algorithms to optimize the policy can lead to performance improvements.
 However, the theoretical justification of performance improvement for such intuition has been less explored.
-For later use, we denote $Z_{mutual}(\cdot|x)$ as the optimal solution for maximizing $I(Z;M)$.
-Whether $Z_{mutual}(\cdot|x)$ is equivalent to $Z^{*}(\cdot|x)$ is still an open problem.
-Hence, we introduce an extra notation here to avoid confusion.
-Theorem 3.1 ([li2024towards]).
+For later use, we denote $Z(\cdot|x;\phi^{mutual})$ as the optimal solution for maximizing $I(Z;M)$.
+Whether $Z(\cdot|x;\phi^{mutual})$ is equivalent to $Z(\cdot|x;\phi^{*})$ remains an open problem.
+To avoid confusion, we introduce an extra notation here.
+Theorem 3.1 ((Li et al., 2024)).
 Denote $X_{b}$ and $X_{t}$ are the behavior-related $(s,a)$-component and task-related $(s^{\prime},r)$-component of the context $X$, with $X=(X_{b},X_{t})$. We have:
 $I(Z;X_{t}|X_{b})\leq I(Z;M)\leq I(Z;X)$ (2)
-where 1) $\mathcal{L}_{FOCAL}\equiv-I(Z;X)=-I(Z;X_{t}|X_{b})-I(Z;X_{b})$; 2) $\mathcal{L}_{CORRO}\equiv-I(Z;X_{t}|X_{b})$; 3) $\mathcal{L}_{CSRO}\geq(\lambda-1)I(Z;X)-\lambda I(Z;X_{t}|X_{b})$
+where 1) $\mathcal{L}_{FOCAL}\equiv-I(Z;X)=-I(Z;X_{t}|X_{b})-I(Z;X_{b})$; 2) $\mathcal{L}_{CORRO}\equiv-I(Z;X_{t}|X_{b})$; 3) $\mathcal{L}_{CSRO}\geq(\lambda-1)I(Z;X)-\lambda I(Z;X_{t}|X_{b})$, and $\equiv$ denotes equality up to a constant.
 4 Methods
-We first introduce a Lipschitz assumption, a widely used technique in both model-free [song2019efficient, ghosh2022provably] and model-based [ji2022update, zhang2023how] RL frameworks.
+We first introduce a Lipschitz assumption, a widely used technique in both model-free (Song & Sun, 2019; Ghosh et al., 2022) and model-based (Ji et al., 2022; Zhang et al., 2023a) RL frameworks.
 Assumption 4.1.
 The policy function is $L_{z}-$ Lipschitz w.r.t some norm |||||$\cdot$||| | ⋅ | | in the sense that
-$\forall Z_{1}(\cdot|x),Z_{2}(\cdot|x)\in\mathbb{Z},\,\,|\pi(\cdot|s,Z_{1}(\cdot|x);\theta)-\pi(\cdot|s,Z_{2}(\cdot|x);\theta)|\leq L_{z}\cdot|Z_{1}(\cdot|x)-Z_{2}(\cdot|x)|$ (3)
+$\forall Z(\cdot|x;\phi_{1}),Z(\cdot|x;\phi_{2})\in\mathbb{Z},\,\,|\pi(\cdot|s,Z(\cdot|x;\phi_{1});\theta)-\pi(\cdot|s,Z(\cdot|x;\phi_{2});\theta)|\leq L_{z}\cdot|Z(\cdot|x;\phi_{1})-Z(\cdot|x;\phi_{2})|$ (3)
+The proofs for all the following sections are provided in Appendix Section 8.2.
 4.1 A Performance Improvement Perspective Towards Prior Works
 Our goal is to arrive at an optimization objective for COMRL to get the performance improvement guarantee.
-Motivated by the return discrepancy scheme [janner2019trust] in the model-based RL field, we can similarly build a tractable lower bound for $J^{*}(\theta)$.
+Motivated by the return discrepancy scheme (Janner et al., 2019) in the model-based RL field, we can similarly build a tractable lower bound for $J^{*}(\theta)$.
 Definition 4.2 (Return discrepancy in COMRL).
@@
 $J^{*}(\theta)-J(\theta)\geq-|J(\theta)-J^{*}(\theta)|$ (4)
-where $J(\theta)$ is the expected return of the policy $\pi$ among the expectation of tasks conditioning on the learned task embedding $Z(\cdot|x)$ as $J(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]]$.
+where $J(\theta)=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x;\phi);\theta)}}[R_{m}(s,a)]]$ is the expected return of the policy $\pi$ conditioning on the learned task representation $Z(\cdot|x;\phi)$.
 From Definition 4.2, we know that if the estimation error $|J^{*}(\theta)-J(\theta)|$ can be upper-bounded, the lower bound for $J^{*}(\theta)$ can be established simultaneously.
@@
 (4) can be interpreted as a unified training objective for both the context encoder and the policy.
-Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the holistic optimization process can be seen as a two-stage alternating framework with the first stage updating the context encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage updating the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$ as $J^{*}(\theta)\geq J(\theta)-|J^{*}(\theta)-J(\theta)|$.
+Specifically, we can alternate the optimization target to lift the lower bound for $J^{*}(\theta)$, where the whole optimization process can be seen as a two-stage alternating framework with the first stage updating the context encoder to minimize $|J^{*}(\theta)-J(\theta)|$ and the second stage updating the policy to maximize $J(\theta)-|J^{*}(\theta)-J(\theta)|$ as $J^{*}(\theta)\geq J(\theta)-|J^{*}(\theta)-J(\theta)|$.
 Hence, the performance improvement guarantee can be achieved.
-The following theorem induces a tractable upper bound for $|J^{*}(\theta)-J(\theta)|$, connecting the bridge between the intuition of previous works and the theoretical justification concerning the performance improvement guarantee.
+The following theorem induces a tractable upper bound for $|J^{*}(\theta)-J(\theta)|$, bridging the gap between the intuition of previous works and the theoretical justification concerning the performance improvement guarantee.
 Theorem 4.3 (Return bound in COMRL).
 Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
-$|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (5)
-Proof: See Section 8.2 in the Appendix. ∎
+$|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{mutual})|+|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|)$ (5)
 By directly unrolling the Theorem 4.3, we can establish the lower bound of $J^{*}(\theta)$ as:
-$J^{*}(\theta)\geq J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (6)
-Note that $|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$ is a constant w.r.t the learned task representation $Z(\cdot|x)$.
-Hence, this term can be ignored when optimizing $Z(\cdot|x)$.
-Admittedly, there may exist a smarter algorithm to further reduce the gap of $|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$, but this is not our main focus and we leave this to future work.
+$J^{*}(\theta)\geq J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{mutual})|+|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|)$ (6)
+Note that $|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|$ is a constant w.r.t the learned task representation $Z(\cdot|x;\phi)$.
+Hence, this term can be ignored when optimizing $Z(\cdot|x;\phi)$.
+Admittedly, there may exist a smarter algorithm to further reduce the gap of $|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|$, but this is not our main focus and we leave this to future work.
 Theorem 4.3 and Eq.
 (6) indicate the following two principles for the context encoder and the policy.
-For the context encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$.
-For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)\equiv J(\theta)$, as the task representation related terms are constants for policy optimization.
-To maximize $J(\theta)$, we can adopt standard offline RL algorithms [wu2019behavior, fujimoto2021minimalist, kumar2020conservative] similar to [yang2022unified].
-Recall that previous COMRL works adopt alternatively training the context encoder by maximizing $I(Z;M)$, which can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x)-Z_{mutual}(\cdot|x)|$ and training the policy by conditioning on the learned task representation as well as applying the standard offline RL algorithms.
-Therefore, we argue that previous COMRL works can be interpreted by this framework, thus providing an explanation concerning performance improvement guarantee.
-However, this optimization framework only considers the discrepancy between $Z(\cdot|x)$ and $Z_{mutual}(\cdot|x)$, without considering the variation of $Z(\cdot|x)$.
+For the context encoder learning, we should minimize the return bound, namely minimizing $\mathbb{E}_{m,x}|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{mutual})|$.
+For the policy learning, we should lift the lower bound of $J^{*}(\theta)$, namely maximizing $J(\theta)-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{mutual})|+|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|)\equiv J(\theta)$, as the task representation related terms are constants for policy optimization.
+To maximize $J(\theta)$, we can adopt standard offline RL algorithms (Wu et al., 2019; Kumar et al., 2020; Fujimoto & Gu, 2021) similar to (Yang et al., 2022).
+Recall that previous COMRL works adopt alternatingly training the context encoder by maximizing $I(Z;M)$, which can be approximately seen as minimizing $\mathbb{E}_{m,x}|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{mutual})|$ and training the policy by conditioning on the learned task representation as well as applying the standard offline RL algorithms.
+Therefore, we argue that our theoretical analyses provide an explanation for the performance improvement guarantee of previous COMRL works.
+However, this optimization framework only considers the discrepancy between $Z(\cdot|x;\phi)$ and $Z(\cdot|x;\phi^{mutual})$, without considering the variation of $Z(\cdot|x;\phi)$.
 In the next section, we will show that this may violate the monotonicity of the performance improvements.
@@
 In this section, we aim to demonstrate that as the previous optimization framework in Section 4.1 doesn’t model the variation of task representation explicitly, this framework would provide an insufficient condition for monotonic performance improvement.
-We first show conditions for monotonic performance improvement of this framework.
+We first show conditions for monotonic performance improvement of the previous framework.
 Corollary 4.4 (Monotonic performance improvement condition for previous COMRL works).
 When meeting Assumption 4.1, the condition for monotonic performance improvement of previous COMRL works is:
-$\epsilon^{*}_{12}\triangleq J^{*}(\theta_{2})-J^{*}(\theta_{1})\geq\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z^{*}(\cdot|x)|)$ (7)
-Proof: See Section 8.2 in the Appendix. ∎
-As shown in Corollary 4.4, the monotonic performance improvement can be guaranteed with only better approximation to $Z^{*}(\cdot|x)$.
+$\epsilon^{*}_{12}\triangleq J^{*}(\theta_{2})-J^{*}(\theta_{1})\geq\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{*})|)$ (7)
+As shown in Corollary 4.4, $Z(\cdot|x;\phi)$ should be close to $Z(\cdot|x;\phi^{*})$ such that the lower bound is small enough for finding a policy to achieve monotonic performance improvement.
 Next, we will introduce the performance difference bound framework to model variation of task representation.
 Definition 4.5 (Performance difference bound in COMRL).
-Denote $J^{1}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the context encoder and the policy.
-Similarly, denote $J^{2}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the context encoder and the policy.
+For an alternating update process, we denote $J^{1}(\theta_{1})=\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x;\phi_{1});\theta_{1})}}[R_{m}(s,a)]$ as the expected return of the policy $\pi(\cdot|s,Z(\cdot|x;\phi_{1});\theta_{1})$ before update of the context encoder and the policy.
+Similarly, denote $J^{2}(\theta_{2})=\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x;\phi_{2});\theta_{2})}}[R_{m}(s,a)]$ as the expected return of the policy $\pi(\cdot|s,Z(\cdot|x;\phi_{2});\theta_{2})$ after update of the context encoder and the policy.
 The performance difference bound in the COMRL setting can be defined as
@@
 when $C$ is non-negative, the algorithm allows a monotonic performance improvement.
-According to Definition 4.5, we need to find the positive $C$ to improve the performance monotonically. To achieve this, we can derive the lower bound of the performance difference.
+According to Definition 4.5, we need to find a positive $C$ to improve the performance monotonically. To achieve this, we can derive the lower bound of the performance difference.
 Theorem 4.6 (Lower bound of performance difference in COMRL).
 Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1, the lower bound of the performance difference in COMRL can be formalized as:
-$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (9)
-Proof: See Section 8.2 in the Appendix. ∎
-For the need of the monotonic performance improvement guarantee, we further let the right-hand side of Eq. (32) be positive. Then, we would have the following condition to achieve monotonicity.
-$\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]\geq 0$ (10)
+$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi^{*})|+|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi_{1})|]$ (9)
+For achieving the monotonic performance improvement guarantee, we further let the right-hand side of Eq. (9) be positive. Then, we would have the following condition to achieve monotonicity.
+$\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi^{*})|+|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi_{1})|]\geq 0$ (10)
 Compared to the condition in Corollary 4.4, Eq.
-10 presents an additional term $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ to achieve monotonic performance improvement, which corresponds exactly to the previous ignored impacts stemming from the variation of the task representation.
-To achieve monotonic performance improvements, the optimization of task representation should consider not only the approximation to $Z^{*}(\cdot|x)$, but also the magnitude of the update.
-For example, if we assume that $Z_{1}(\cdot|x)$ is trained from scratch and the update process from $Z_{1}(\cdot|x)$ to $Z_{2}(\cdot|x)$ brings $Z_{2}(\cdot|x)$ close to $Z^{*}(\cdot|x)$, then with the condition in Corollary 4.4, the monotonic performance improvement can be easily achieved with small $\epsilon^{*}_{12}$.
+10 presents an additional term $|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi_{1})|$ to achieve monotonic performance improvement, which corresponds exactly to the previous ignored impacts stemming from the variation of the task representation.
+To achieve monotonic performance improvements, the optimization of task representation should consider not only the approximation to $Z(\cdot|x;\phi^{*})$, but also the magnitude of the update.
+For example, if we assume that $Z(\cdot|x;\phi_{1})$ is trained from scratch and the update process from $Z(\cdot|x;\phi_{1})$ to $Z(\cdot|x;\phi_{2})$ brings $Z(\cdot|x;\phi_{2})$ close to $Z(\cdot|x;\phi^{*})$, then with the condition in Corollary 4.4, the monotonic performance improvement can be easily achieved with small $\epsilon^{*}_{12}$.
 However, for the condition in Eq.
-(10), small $\epsilon^{*}_{12}$ may cause the violation of monotonicity as $|Z_{1}(\cdot|x)-Z_{2}(\cdot|x)|$ remains large.
-To highlight the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, we name this issue Task Representation Shift.
+(10), small $\epsilon^{*}_{12}$ may cause the violation of monotonicity as $|Z(\cdot|x;\phi_{1})-Z(\cdot|x;\phi_{2})|$ remains large.
+To highlight the impacts of $|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi_{1})|$, we name this issue Task Representation Shift.
 Under some mild assumptions, we can conclude that it is possible to achieve monotonic performance improvement with sufficient policy improvement $\epsilon^{*}_{12}$ on the condition of appropriate encoder updates, as shown in Theorem 4.10.
@@
 Assumption 4.9.
-There exists an $\alpha$ for any given $Z_{1}(\cdot|x)$, $|Z_{2}(\cdot|x)-\tilde{Z}_{2}(\cdot|x)|^{2}\leq\frac{\alpha}{b}$, where $Z_{2}(\cdot|x)$ denotes the context encoder updated by maximizing $I(Z;M)$ with the data size $b$ randomly sampled from the training dataset based on $Z_{1}(\cdot|x)$ and $\tilde{Z}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $b$ i.i.d samples from $Z^{*}(\cdot|x)$ based on $Z_{1}(\cdot|x)$  [^1] We set discrepancy to be subject to an inverse proportional function is because as the training samples increase, $Z_{2}$ and $\tilde{Z}_{2}$ should become similar..
+There exists an $\alpha$ for any given $Z(\cdot|x;\phi_{1})$, $|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\tilde{\phi}_{2})|^{2}\leq\frac{\alpha}{b}$, where $Z(\cdot|x;\phi_{2})$ denotes the context encoder updated by maximizing $I(Z;M)$ with the data size $b$ randomly sampled from the training dataset based on $Z(\cdot|x;\phi_{1})$ and $Z(\cdot|x;\tilde{\phi}_{2})$ denotes the context encoder updated by fitting the empirical distribution on $b$ i.i.d samples from $Z(\cdot|x;\phi^{*})$ based on $Z(\cdot|x;\phi_{1})$  [^2] For more details to justify these assumptions, please refer to Appendix 8.3..
 Theorem 4.10 (Monotonic performance improvement guarantee on training process).
-Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex.
+Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ and $|Z|$ as the cardinality of the task representation space.
 Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent.
 When meeting Assumption 4.1, 4.7, 4.8 and 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder via maximizing $I(Z;M)$ from at least extra $k$ samples, where:
-$k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (11)
+$k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{|Z|}-2}{\xi}}+\sqrt{\alpha})^{2}$ (11)
 Here, $\xi\in[0,1]$ is a constant.
-Proof: See Section 8.2 in the Appendix. ∎
 Theorem 4.10 shows the connection between the needed update data size $k$ and the performance improvement brought only by the policy update $\epsilon^{*}_{12}$.
@@
 Algorithm 1 General Algorithmic Framework Towards Reining In The Task Representation Shift
-0:  Offline training datasets $\bm{X}$, initialized policy $\pi_{\theta}$, context encoder $q_{\phi}$ and given task representation shift threshold $\beta$
+0:  Offline training datasets $\bm{X}$, initialized policy $\pi_{\theta}$, context encoder $Z_{\phi}$, given task representation shift threshold $\beta$ and given training batch size for context encoder $N_{\text{bs}}$.
 1:  for iter in alternating iterations do
 2:     // Update the context encoder
-3:     if Conditions for updating the context encoder are met then
-4:        while Accumulated task representation shift is less than $\beta$ do
-5:           Sample context from $\bm{X}$
-6:           Obtain task representations from $q_{\phi}$ with inputting the sampled context
-7:           Compute $\mathcal{L}_{\text{encoder}}$ concerning maximizing $I(Z;M)$
-8:           Update $\phi$ to minimize $\mathcal{L}_{\text{encoder}}$
-9:        end while
-10:     end if
-11:     // Update the policy
-12:     Detach the task representations
-13:     Sample training data from $\bm{X}$
-14:     Compute $\mathcal{L}_{\text{policy}}$ via standard offline RL algorithms
-15:     Update $\theta$ to minimize $\mathcal{L}_{\text{policy}}$
-16:  end for
+3:     Estimate the conditions (e.g.use Eq. (11) to approximate $k$)
+4:     if Conditions for updating the context encoder are met (e.g. $k<=N_{\text{bs}}$) then
+5:        while Accumulated task representation shift is less than $\beta$ do
+6:           Sample context from $\bm{X}$
+7:           Obtain task representations from $Z_{\phi}$ with inputting the sampled context
+8:           Compute $\mathcal{L}_{\text{encoder}}$ concerning maximizing $I(Z;M)$
+9:           Update $\phi$ to minimize $\mathcal{L}_{\text{encoder}}$
+10:        end while
+11:     end if
+12:     // Update the policy
+13:     Detach the task representations
+14:     Sample training data from $\bm{X}$
+15:     Compute $\mathcal{L}_{\text{policy}}$ via standard offline RL algorithms
+16:     Update $\theta$ to minimize $\mathcal{L}_{\text{policy}}$
+17:  end for
 4.3 Practical Implementation
-To show the impacts of the task representation shift issue, we use different settings to adjust the context encoder.
-We denote the given batch size as $bs$ and the accumulation steps of task representation shift as $acc$.
-Our settings include 1) $k=2\times bs,acc=1$, 2) $k=3\times bs,acc=1$, 3) $k=1\times bs,acc=2$, and 4) $k=1\times bs,acc=3$.
-All settings as well as the original setting, namely $k=1\times bs,acc=1$ are performed for 8 different random seeds.
-With respect to the specific context encoder learning algorithms concerning maximizing $I(Z;M)$, we choose 3 representative algorithms that are also widely adopted in previous COMRL endeavors.
-The reconstruction-based algorithm, passing the state, action, and task representation through the decoder to reconstruct the next state and the reward signal, is applied in [zintgraf2019varibad, dorfman2021offline, li2024towards] and is proven to be the lower bound of $I(Z;M)$.
-The contrastive-based algorithm, using the task representation to compute distance metric learning loss, is applied in [li2020focal, gao2024context, li2024towards] and is proven to be the upper bound of $I(Z;M)$ as well as contain the lower bound of $I(Z;M)$.
-The cross-entropy-based algorithm is a straightforward discrete approximation to $I(Z;M)$ and is proposed by our work.
-Specifically, after extracting from the context encoder, the task representation would be passed through a fully connected layer to get the probabilities for each task.
-Then, the context encoder is trained by the supervision of the task label and back-propagation of the cross-entropy loss.
-The implementation details of these algorithms are shown in Appendix Section 8.3.
-With respect to the policy learning algorithm, to maintain the consistency of previous COMRL works, we directly adopt BRAC [wu2019behavior] to train the policy.
+As outlined in Algorithm 1, the way to rein in the task representation shift can be seen as two aspects, namely 1) when to update the context encoder determined by $k$ and 2) how many times to update the context encoder determined by $\beta$.
+To cover these two aspects, we introduce two parameters $N_{\text{k}}$ and $N_{\text{acc}}$.
+Here, $N_{\text{k}}=n$ denotes that the context encoder needs to be updated every $n$ updates of the policy, and $N_{\text{acc}}=n$ denotes when the context encoder needs to be updated, it is updated $n$ times.
+Our settings include 1) $N_{\text{k}}=2,N_{\text{acc}}=1$, 2) $N_{\text{k}}=3,N_{\text{acc}}=1$, 3) $N_{\text{k}}=1,N_{\text{acc}}=2$, and 4) $N_{\text{k}}=1,N_{\text{acc}}=3$.
+All settings as well as the original setting, namely $N_{\text{k}}=1,N_{\text{acc}}=1$, are performed for 8 different random seeds.
+With respect to the specific context encoder learning algorithms concerning maximizing $I(Z;M)$, we choose three representative algorithms that have been widely adopted in previous COMRL endeavors.
+The details of these algorithms are shown in Appendix Section 8.4.
+Contrastive-based is applied in (Li et al., 2020; Gao et al., 2024; Li et al., 2024) and it is proven to be the upper bound of $I(Z;M)$.
+Reconstruction-based is applied in (Zintgraf et al., 2019; Dorfman et al., 2021; Li et al., 2024) and it is proven to be the lower bound of $I(Z;M)$, which is equivalent to CORRO (Yuan & Lu, 2022) under the offline setting.
+Cross-entropy-based is proposed in our work and it is a direct approximation w.r.t $I(Z;M)$.
+With respect to the policy learning algorithm, to maintain the consistency of previous COMRL works, we directly adopt BRAC (Wu et al., 2019) to train the policy.
 5 Experiment
@@
 5.1 Environments Settings
-We adopt MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta] benchmarks to evaluate the algorithms.
-Following the protocol of UNICORN  [li2024towards], we randomly sample 20 training tasks and 20 testing tasks from the task distribution.
-We train the SAC [haarnoja2018soft] agent from scratch for each task and use the collected replay buffer as the offline dataset.
-During the meta-testing phase, we adopt a fully offline setting in a few-shot manner like [li2020focal] that randomly samples a trajectory from the dataset as the context to obtain the task representation and then passes it to the meta-policy to complete the testing in the true environments.
+We adopt MuJoCo (Todorov et al., 2012) and MetaWorld (Yu et al., 2020) benchmarks to evaluate the algorithms.
+Following the protocol of UNICORN  (Li et al., 2024), we randomly sample 20 training tasks and 20 testing tasks from the task distribution.
+We train the SAC (Haarnoja et al., 2018) agent from scratch for each task and use the collected replay buffer as the offline dataset.
+During the meta-testing phase, we adopt a fully offline setting in a few-shot manner like (Li et al., 2020) that randomly samples a trajectory from the dataset as the context to obtain the task representation and then passes it to the meta-policy to complete the testing in the true environments.
 5.2 MAIN RESULTS
-Figure 2: Testing returns of different settings to rein in the task representation shift against the original setting concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$.
+Figure 2: Testing returns of different settings to rein in the task representation shift.
 Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
-Figure 2 illustrates the learning curves of our different settings to rein in the task representation shift against the original setting on 20 testing tasks for each benchmark concerning 3 widely adopted objectives for maximizing $I(Z;M)$.
-We find that these specific objectives hold unanimous performance demonstration on all benchmarks that the previous training framework, namely $k=1\times bs,acc=1$, cannot provide the best asymptotic performance.
-These objectives focus solely on the optimization target of the task representation, namely the optimal solution of $I(Z;M)$.
-While they are guaranteed to produce reasonable results, as shown in Section 4.1, ignoring the impacts of task representation shift still hinders the capability of these algorithms to reach their full potential.
-As shown by Figure 2, even with minor changes to the algorithms, the performance improvements are substantial.
-Figure 3: Testing returns of different manners to rein in the task representation shift against the original manner concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$ on different data qualities of Ant-Dir.
+Figure 2 and Table 3 illustrate the performance of our different settings to rein in the task representation shift on 20 testing tasks.
+We find that these specific objectives exhibit clear trends across all benchmarks that the best setting showcases statistically significant performance improvements compared to the worst setting, thereby emphasizing the indispensability of considering the task representation shift issue.
+We also notice that the previous training framework, namely $N_{\text{k}}=1,N_{\text{acc}}=1$ , fails to deliver the best asymptotic performance across all benchmarks.
+Though $N_{\text{k}}=1,N_{\text{acc}}=1$ can be seen as an implicit way to rein in the task representation shift, the lack of explicit control still hinders the capability of these base algorithms to reach their full potential.
+Furthermore, as shown in Table 3, we observe that settings with $N_{\text{k}}>1,N_{\text{acc}}=1$ achieve better performance more frequently than those with $N_{\text{k}}=1,N_{\text{acc}}>1$ w.r.t the original setting $N_{\text{k}}=1,N_{\text{acc}}=1$.
+Based on this observation, we recommend prioritizing the adjustment of $N_{\text{k}}$ when tuning parameters.
+This not only tends to yield better performance but also offers the advantage of reducing training time.
+In contrast, increasing $N_{\text{acc}}$ would increase training time due to the need for multiple context encoder updates within a single alternating step.
+Figure 3: Testing returns of different settings to rein in the task representation shift on different data qualities in Ant-Dir.
 Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
@@
 Depending on the quality of the behavior policies, we denote these datasets as random, medium, and expert respectively.
-Figure 3 demonstrates the learning curves of our different settings to rein in the task representation shift against the original setting concerning 3 widely adopted objectives for maximizing $I(Z;M)$ on Ant-Dir.
+Figure 3 and Table 3 demonstrate the performance of our different settings to rein in the task representation shift on Ant-Dir.
 We find that the contrastive-based algorithm fails on the random dataset.
-Except for this case, the others show consistent results that reining in the task representation shift may lead to better performance improvements.
+Except for this case, the others show consistent results that reining in the task representation shift can lead to better performance improvements.
 Based on these impressive results, we believe that developing a smarter algorithm to control the task representation shift automatically is an appealing direction and we leave this to future work.
@@
 This section seeks to raise some interesting issues derived from our work. We hope these issues can further enable pondering on the significance of the task representation shift.
-Figure 4: Testing returns of the pre-training scheme against training from scratch of cross-entropy, contrastive, and reconstruction based objectives on Ant-Dir.
+Figure 4: Testing returns of the pre-training scheme against training from scratch in Ant-Dir.
 Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
@@
 If we train the context encoder in advance and use this pre-trained context encoder to train the policy directly, can we achieve better performance improvement?
-To answer this question, we conduct the experiment on Ant-Dir and still use the cross-entropy, contrastive, and reconstruction based objectives to pre-train the context encoder.
+To answer this question, we conduct the experiment in Ant-Dir and still use the cross-entropy, contrastive, and reconstruction based objectives to pre-train the context encoder.
 As shown in Figure 4, there is a significant performance gap between pre-training and training from scratch.
@@
 Corollary 6.1 (Monotonic performance improvement condition for pre-training scheme).
-Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, the monotonic performance improvement condition for pre-training scheme is:
-$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (12)
-Proof: See Section 8.2 in the Appendix. ∎
-Due to the implementation error and approximation error, $|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|$ can not equal to 0.
+Denote $Z(\cdot|x;\phi^{pretrain})$ as the task representation distribution after pre-training. When meeting Assumption 4.1, the monotonic performance improvement condition for pre-training scheme is:
+$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z(\cdot|x;\phi^{pretrain})-Z(\cdot|x;\phi^{*})|]\geq 0$ (12)
+While maximizing $I(Z;M)$ can help $Z(\cdot|x;\phi^{pretrain})$ to approach $Z(\cdot|x;\phi^{mutual})$, the sub-optimal gap between $Z(\cdot|x;\phi^{mutual})$ and $Z(\cdot|x;\phi^{*})$ still persists.
+Furthermore, given the inherent approximation error between $Z(\cdot|x;\phi^{pretrain})$ and $Z(\cdot|x;\phi^{mutual})$, the discrepancy between $Z(\cdot|x;\phi^{pretrain})$ and $Z(\cdot|x;\phi^{*})$ maybe large.
 Hence, according to Corollary 6.1, pre-training cannot actually achieve monotonic performance improvements.
 As the task representation cannot be changed under the pre-training scheme, compared to training from scratch, this scheme loses some degrees of freedom to improve performance.
-Additionally, this scheme can be intuitively regarded as degrading to the problem that how to design a pre-training algorithm to better approximate $Z^{*}(\cdot|x)$.
-Notice that there exists the decomposition $|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|\leq|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$.
-Since the cross-entropy based objective is the direct approximation of $I(Z;M)$, holding better performance under the pre-training scheme is also in line with expectation.
-This analysis further enhances the importance of considering the impacts of task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ in the alternating process.
+Additionally, this scheme can be intuitively regarded as degrading to the problem that how to design a pre-training algorithm to better approximate $Z(\cdot|x;\phi^{*})$.
+Notice that there exists the decomposition $|Z(\cdot|x;\phi^{pretrain})-Z(\cdot|x;\phi^{*})|\leq|Z(\cdot|x;\phi^{pretrain})-Z(\cdot|x;\phi^{mutual})|+|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|$.
+Since the cross-entropy-based objective is the direct approximation of $I(Z;M)$, holding better performance under the pre-training scheme is also in line with expectation.
+This analysis further enhances the importance of considering the impacts of task representation shift $|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi_{1})|$ in the alternating process.
 Nevertheless, how to achieve better performance under the pre-training scheme is also an interesting problem.
@@
 6.2 Can The Visualization Of Task Representation Be Strongly Relied Upon to Imply The Asymptotic Performance?
-Previous COMRL endeavors [li2020focal, yuan2022robust, gao2024context] mostly apply t-SNE on the learned task representation to show the differentiation of the tasks.
+Previous COMRL endeavors (Li et al., 2020; Yuan & Lu, 2022; Gao et al., 2024) mostly apply t-SNE on the learned task representation to show the differentiation of the tasks.
 They hold the belief that better performance corresponds to better differentiation of the tasks.
-However, based on our analysis, the visualization result only depends on the task representation at convergence and ignores the task representation variation during the whole optimization process.
+However, based on our analysis, the visualization result only depends on the task representation at convergence and ignores the task representation shift during the whole optimization process.
 According to the illustration shown in Figure 5, less desirable differentiation results can also lead to better performance.
-Hence, it is insufficient to imply performance based on such evaluation principles.
-Nevertheless, we recognize that the visualization result can be seen as an auxiliary metric to assist in determining the task representation.
+Hence, the visualization results may represent the true task distribution but cannot sufficiently imply the final performance.
 7 Conclusion & Limitation
 Finding that the performance improvement guarantee of the COMRL training framework remains under-explored, we constructively provide a theoretical framework to link the training scheme in COMRL with the general RL objective of maximizing the expected return.
-After scrutinizing this framework, we further find it ignores the variation of the task representation, which weakens the condition necessary for monotonic performance improvements.
+After scrutinizing this framework, we further find it ignores the variation of the task representation, which may impair performance improvement.
 Based on this finding, we propose a new issue called task representation shift, refine the condition for monotonicity, and prove that monotonic performance improvements can be achieved with appropriate context encoder updates.
-With minor changes to the previous training framework for reining in the task representation shift, the performance improvements can be substantial.
 We prospect that deeply exploring the role of task representation shift can make a profound difference in the COMRL setting.
@@
 We leave these interesting problems to future work.
-[…]
-8 Appendix
-8.1 Useful Lemmas
-Lemma 8.1 (Return bound.[zhang2023how]).
-Let $R_{max}$ denote the bound of the reward function, $\epsilon_{\pi}$ denote $\max_{s}D_{TV}(\pi_{1}||\pi_{2})$ and $\epsilon^{M_{2}}_{M_{1}}$ denote $\mathbb{E}_{(s,a)\sim d^{\pi_{1}}_{M_{1}}}[D_{TV}(p_{M_{1}}||p_{M_{2}})]$.
-For two arbitrary policies $\pi_{1},\pi_{2}\in\Pi$, the expected return under two arbitrary models $M_{1},M_{2}\in\mathcal{M}$ can be bounded as,
-$|V^{\pi_{2}}_{M_{2}}-V^{\pi_{1}}_{M_{1}}|\leq 2R_{max}(\frac{\epsilon_{\pi}}{(1-\gamma)^{2}}+\frac{\gamma}{(1-\gamma)^{2}}\epsilon^{M_{2}}_{M_{1}})$ (13)
-Proof:
-$\begin{split}|V^{\pi_{2}}_{M_{2}}-V^{\pi_{1}}_{M_{1}}|&=|\sum^{\infty}\limits_{t=0}\gamma^{t}\sum\limits_{s,a}(p^{\pi_{2}}_{t,M_{2}}(s,a)-p^{\pi_{1}}_{t,M_{1}}(s,a))r(s,a)|\\ &\leq R_{max}\sum^{\infty}\limits_{t=0}\gamma^{t}\sum\limits_{s,a}|p^{\pi_{2}}_{t,M_{2}}(s,a)-p^{\pi_{1}}_{t,M_{1}}(s,a)|\\ &=2R_{max}\sum^{\infty}\limits_{t=0}\gamma^{t}D_{TV}(p^{\pi_{1}}_{t,M_{1}}(s,a)||p^{\pi_{2}}_{t,M_{2}}(s,a))\\ \end{split}$ (14)
-According to Theorem 2 (Return Bound) in [zhang2023how],
-$\sum^{\infty}\limits_{t=0}\gamma^{t}D_{TV}(p^{\pi_{1}}_{t,M_{1}}(s,a)||p^{\pi_{2}}_{t,M_{2}}(s,a))\leq(\frac{\epsilon_{\pi}}{(1-\gamma)^{2}}+\frac{\gamma}{(1-\gamma)^{2}}\epsilon^{M_{2}}_{M_{1}})$ (15)
-We bring this result back and can draw the conclusion.
-Lemma 8.2 (Inequility for $L_{1}$ deviation of the empirical distribution.[weissman2003inequalities]).
-Let $P$ be a probability distribution on the set $\mathcal{A}=\{1,...,a\}$.
-For a sequence of samples $x_{1},...,x_{m}\sim P$, let $\hat{P}$ be the empirical probability distribution on $\mathcal{A}$ defined by $\hat{P}(j)=\frac{1}{m}\sum_{i=1}^{m}\mathbb{I}(x_{i}=j)$.
-The $L_{1}$-deviation of the true distribution $P$ and the empirical distribution $\hat{P}$ over $\hat{A}$ from $m$ independent identically samples is bounded by,
-$Pr(|P-\hat{P}|\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp{(-m\epsilon^{2}/2)}$ (16)
-Proof:
-According to the inequality for $L_{1}$ deviation of the empirical distribution [weissman2003inequalities], they conclude that $Pr(|P-\hat{P}|_{1}\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp(-m\phi(\pi_{P})\epsilon^{2}/4)$, where $\phi(p)=\frac{1}{1-2p}\log\frac{1-p}{p}$ and $\pi_{P}=\max_{A\in\mathcal{A}}\min(P(A),1-P(A))$.
-The paper [weissman2003inequalities] point out that $0\leq\pi_{P}\leq 1/2,\forall P$.
-Based on this, we can derive the function $\phi(p)$ and find that $\phi(p)$ is monotonically decreasing on $(0,1/2)$.
-Therefore, we can get $\phi(p)\geq\phi(1/2)=2$.
-Then, we bring this result back and can conclude that $Pr(|P-\hat{P}|_{1}\geq\epsilon)\leq(2^{|\mathcal{A}|}-2)\exp(-m\epsilon^{2}/2)$.
-8.2 Missing Proofs
-Theorem 8.3 (Return bound in COMRL).
-Assume the reward function is upper-bounded by $R_{max}$. For an arbitrary policy parameter $\theta$, when meeting Assumption 4.1, the return bound in COMRL can be formalized as:
-$|J^{*}(\theta)-J(\theta)|\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (17)
-Proof:
-$|J(\theta)-J^{*}(\theta)|$
-$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]]-\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (18)
-$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (19)
-$=|\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]$
-$+\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]]|$ (20)
-$\leq\mathbb{E}_{m,x}[|\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]|$
-$+|\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}}[R_{m}(s,a)]|]$ (21)
-$=\mathbb{E}_{m,x}[|V^{\pi(\cdot|s,Z(\cdot|x);\theta)}_{m}-V^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{m}|+|V^{\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)}_{m}-V^{\pi(\cdot|s,Z^{*}(\cdot|x);\theta)}_{m}|]$ (22)
-$\leq\mathbb{E}_{m,x}[\frac{2R_{max}}{(1-\gamma)^{2}}\max_{s}D_{TV}(\pi(\cdot|s,Z(\cdot|x);\theta)||\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta))$
-$+\frac{2R_{max}}{(1-\gamma)^{2}}\max_{s}D_{TV}(\pi(\cdot|s,Z_{mutual}(\cdot|x);\theta)||\pi(\cdot|s,Z^{*}(\cdot|x);\theta))]$ (23)
-$\leq\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|)$ (24)
-Eq. (23) is the result of directly applying Lemma 8.1 and Eq. (24) is the result of directly applying Assumption 4.1. ∎
-Corollary 8.4 (Monotonic performance improvement condition for previous COMRL works).
-When meeting Assumption 4.1, the condition for monotonic performance improvement of previous COMRL works is:
-$\epsilon^{*}_{12}\triangleq J^{*}(\theta_{2})-J^{*}(\theta_{1})\geq\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z^{*}(\cdot|x)|)$ (25)
-Proof:
-$J(\theta_{2})-J(\theta_{1})=J(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J(\theta_{1})$ (26)
-Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1 and Assumption 4.1 into the above formulation, we can get:
-$J(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J(\theta_{1})$ (27)
-$\geq\mathbb{E}_{m,x}[-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z(\cdot|x);\theta_{2})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{2}))]+\epsilon^{*}_{12}$ (28)
-$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z(\cdot|x);\theta_{1})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{1}))]$ (29)
-$\geq\mathbb{E}_{m,x}[-2\kappa L_{z}|Z(\cdot|x)-Z^{*}(\cdot|x)|+\epsilon^{*}_{12}]$ (30)
-To get the monotonic performance improvement, we need Eq. (30) to be larger than 0. Hence, we can get:
-$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (31)
-Theorem 8.5 (Lower bound of performance difference in COMRL).
-Assume the reward function is upper-bounded by $R_{max}$. When meeting Assumption 4.1, the lower bound of the performance difference in COMRL can be formalized as:
-$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (32)
-Proof: We can introduce the following decomposition
-$J^{2}(\theta_{2})-J^{1}(\theta_{1})$ (33)
-$=\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]$ (34)
-$=\mathbb{E}_{m,x}[\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]]$ (35)
-Denote $\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{i}(\cdot|x);\theta_{j})}}[R_{m}(s,a)]]$ as $G^{\theta_{j}}_{Z_{i}}$, then we can get the following derivation.
-Here, we only consider the value within the brackets.
-$=\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})}}[R_{m}(s,a)]-\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})}}[R_{m}(s,a)]$ (36)
-$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (37)
-$=G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z^{*}}+G^{\theta_{2}}_{Z^{*}}-G^{\theta_{1}}_{Z^{*}}+G^{\theta_{1}}_{Z^{*}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (38)
-Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1 and Assumption 4.1 into the above formulation, we can get:
-$G^{\theta_{2}}_{Z_{2}}-G^{\theta_{2}}_{Z^{*}}+G^{\theta_{2}}_{Z^{*}}-G^{\theta_{1}}_{Z^{*}}+G^{\theta_{1}}_{Z^{*}}-G^{\theta_{1}}_{Z_{2}}+G^{\theta_{1}}_{Z_{2}}-G^{\theta_{1}}_{Z_{1}}$ (39)
-$\geq-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{2}))+\epsilon^{*}_{12}$
-$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1}))$
-$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1}))$ (40)
-$\geq-\kappa L_{z}|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+\epsilon^{*}_{12}$
-$-\kappa L_{z}|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|-\kappa L_{z}|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ (41)
-Simplifying Eq. (41) we can get:
-$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (42)
-Theorem 8.6 (Monotonic performance improvement guarantee on training process).
-Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex.
-Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent.
-When meeting Assumption 4.1, 4.7, 4.8 and 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder via maximizing $I(Z;M)$ from at least extra $k$ samples, where:
-$k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (43)
-Here, $\xi\in[0,1]$ is a constant.
-Proof:
-Let $Z_{1}$ have already been trained by maximizing $I(Z;M)$ to some extent.
-Given that to get $Z_{2}$, we need to train $Z_{1}$ by extra $k$ samples from the training dataset by maximizing $I(Z;M)$ to get the monotonic performance improvement guarantee.
-We begin with the following inequality:
-$|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|\leq|Z_{2}(\cdot|x)-\tilde{Z}_{2}(\cdot|x)|+|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|$ (44)
-where $\tilde{Z}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $k$ i.i.d samples from $Z^{*}(\cdot|x)$ based on $Z_{1}(\cdot|x)$.
-According to Assumption 4.9, we have:
-$|Z_{2}(\cdot|x)-\tilde{Z}_{2}(\cdot|x)|\leq\sqrt{\frac{\alpha}{k}}$ (45)
-Now, we move our focus on solving $|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|$.
-By applying Lemma 8.2, the $L_{1}$ deviation of the empirical distribution $\tilde{Z}_{2}(\cdot|x)$ and true $Z^{*}(\cdot|x)$ over $vol(\mathbb{Z})$ is bounded by:
-$Pr(|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|\geq\epsilon)\leq(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (46)
-$Pr(|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|<=\epsilon)\geq 1-(2^{vol(\mathcal{\mathbb{Z}})}-2)\exp{(-k\epsilon^{2}/2)}$ (47)
-Then for a fixed $x$, with probability greater than $1-\xi$, we have:
-$|\tilde{Z}_{2}(\cdot|x)-Z^{*}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}$ (48)
-Hence, we can get:
-$|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|\leq\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}$ (49)
-Let $\epsilon=\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$, and recall that $\kappa$ denotes $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$, then we can get the $k$ as:
-$\sqrt{\frac{2}{k}\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\frac{\alpha}{k}}\leq\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ (50)
-$k\geq\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (51)
-Since we need the least number of samples that update the context encoder, we take $k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$. ∎
-Corollary 8.7 (Monotonic performance improvement condition for pre-training scheme).
-Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, the monotonic performance improvement condition for pre-training scheme is:
-$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (52)
-Proof: Denote $J^{pretrain}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the policy.
-Similarly, denote $J^{pretrain}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the policy.
-$J^{pretrain}(\theta_{2})-J^{pretrain}(\theta_{1})$ (53)
-$=J^{pretrain}(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J^{pretrain}(\theta_{1})$ (54)
-Denote $\frac{2R_{max}}{(1-\gamma)^{2}}$ as $\kappa$. Taking Lemma 8.1 and Assumption 4.1 into the above formulation, we can get:
-$J^{pretrain}(\theta_{2})-J^{*}(\theta_{2})+J^{*}(\theta_{2})-J^{*}(\theta_{1})+J^{*}(\theta_{1})-J^{pretrain}(\theta_{1})$ (55)
-$\geq\mathbb{E}_{m,x}[-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{2})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{2}))]+\epsilon^{*}_{12}$ (56)
-$-\kappa\max_{s}D_{TV}(\pi(\cdot|s,Z_{pretrain}(\cdot|x);\theta_{1})||\pi(\cdot|s,Z^{*}(\cdot|x);\theta_{1}))]$ (57)
-$\geq\mathbb{E}_{m,x}[-2\kappa L_{z}|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|+\epsilon^{*}_{12}]$ (58)
-To get the monotonic performance improvement, we need Eq. (58) to be larger than 0. Hence, we can get:
-$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (59)
-8.3 Implementation Details
-In this section, we report the implementation details of all the objectives in our main paper.
-  • Contrastive-based [li2020focal]. We adopt the open-source code of FOCAL [^2] https://github.com/FOCAL-ICLR/FOCAL-ICLR/.
-  • Reconstruction-based [li2024towards]. We obtain the source code of UNICORN from the authors and adopt UNICORN-0 as the reconstruction-based objective training framework.
-  • Cross-entropy-based. Built upon the code base of FOCAL, cross-entropy-based only replaces the distance metric learning loss with cross-entropy loss.
-To maintain fairness, we make sure all benchmark-irrelevant parameters are consistent, as shown in Table 1.
-Table 1: Benchmark-irrelevant parameters setting in the training process.
-training tasks 20
-testing tasks 20
-task training batch size 16
-rl batch size 256
-context size 1 trajectory
-actor-network size [256, 256]
-critic-network size [256, 256]
-task encoder network size [64, 64]
-learning rate 3e-4
-Table 2: Benchmark-relevant parameters setting in the training process.
-Configurations Reach Ant-Dir Button-Press Dial-Turn
-dataset size 3e5 9e4 3e5 3e5
-task representation dimension 5 5 5 5
-For each environment, we make the benchmark-relevant parameters the same for each algorithm, as shown in Table 2.
+Acknowledgments
+The work is supported by the National Key Research and Development Program of China (No.
+2020YFA0711402) and by the Young Scientists Fund of National Natural Science Foundation of China (No.
+62406295).